### v1

# Class-Discriminative Attention Maps for Vision Transformers

Abstract
Interpretability methods are critical components for examining and exploring deep neural networks (DNN), as well as increasing our understanding of and trust in them. Vision transformers (ViT), which can be trained to state-of-the-art performance with a self-supervised learning (SSL) training method [1], provide built-in attention maps (AM). While AMs can provide high-quality semantic segmentation of input images, they do not account for any signal coming from a downstream classifier. We introduce class-discriminative attention maps (CDAM), a novel post-hoc explanation method that is highly sensitive to the target class. Our method essentially scales attention scores by how relevant the corresponding tokens are for the predictions of a classifier head. Alternative to classifier outputs, CDAM can also explain a user-defined concept by targeting similarity measures in the latent space of the ViT. This allows for explanations of arbitrary concepts, defined by the user through a few sample images. We investigate the operating characteristics of CDAM in comparison with relevance propagation (RP) [2] and token ablation maps (TAM), an alternative to pixel occlusion methods. CDAM is highly class-discriminative and semantically relevant, while providing implicit regularization of relevance scores.
PyTorch implementation: https://github.com/lenbrocki/CDAM.
Web live demo: https://cdam.informatism.com/.
Figure 1: Extending attention maps (AM), the proposed class-discriminative attention maps (CDAM) quantify and visualize input features that are relevant for the target class in Vision Transformers (ViT). We trained a linear classifier on the ViT-S/8 trained with DINO as a backbone [1] and visualize random samples with CDAM for different target classes. For details, refer to Section 3.1. Orange and blue colors correspond to positive and negative values, respectively.
1 Introduction
The transformer architecture, which was originally developed for natural language processing [3], has been developed into vision transformers (ViT) that demonstrate comparable or even better performance than convolutional neural networks (CNN) on many computer vision tasks [4, 5]. Additionally, a powerful feature of transformers is that the attention mechanism provides a built-in explanation of the model’s inner workings, if attention is understood as a weighing of the relevance of individual tokens [6, 7, 8, 9]. While attention maps (AM) provides intuitive interpretability, it fails to be discriminative with respect to a target class. Recognizing the potential of AM for explaining a ViT’s inner workings, we propose class-discriminative attention maps (CDAM), a gradient-based extension that makes them sensitive to class-specific signals coming from a downstream classifier.
Recently, it has been demonstrated that ViT models trained with DINO, a self-supervised learning (SSL) training method [1], produce AMs that provide high-quality semantic segmentations of input images. These AMs clearly distinguish objects of interest from the background with very little noise, with different attention heads often focusing on semantically distinct regions [1]. AMs are the visualization of the self-attention of the [CLS] token query in the final attention layer. However, one shortcoming of AMs is that they purely operate on the level of the last attention layer and therefore do not take into account any information coming from a downstream task, such as a classifier head.
The proposed CDAMs rely on the calculation of gradients of the class score, i.e. the logit value of a target class, with respect to the token activations in the final transformer block (Fig. 2(a)). We show through explicit calculation and empirical observations that CDAM scales the attention by how relevant the corresponding tokens are for the model’s decision. CDAM therefore inherits the desirable properties of AMs, but additionally indicates the (counter-)evidence for a target class that the ViT relies on for its predictions (Fig. 1). We demonstrate that it is class-discriminative in the sense that the heat maps clearly distinguish the object representing the target class from the background and, importantly, from objects that belong to other classes. CDAMs are also highly sensitive to the target class which, all together, makes it a useful tool for explaining the decision-making process of classifiers with a ViT backbone.
In addition to the classifier-based explanations, we present a concept-based approach that replaces the class score with a similarity measure in the latent space of the ViT, where a concept is defined through a number of example images. Averaging the latent representations of the example images yields a concept embedding, which is analogous to a concept vector in variational autoencoders (VAE) and related generative models [10, 11]. In brief, we use CDAM to identify image patches that drive the (dis-)similarity of the concept embedding and the latent representation of a given image. This allows obtaining a heat map for the target concept the model has never been explicitly trained on, which is of particular interest for self-supervised models. Both supervised and semi-supervised variants of CDAM are compared to token ablation maps (TAMs) and relevance propagation (RP) [2]. Overall, our results suggest that CDAMs exhibit semantic coherency, implicit regularization, and strong discrimination with respect to a target class/concept.
2 Related Works
Interpretability of ViT, and more broadly deep neural networks (DNN), is an active area of research, since understanding their decision-making process would help improve the model, identify weaknesses and biases, and ultimately increase human trust in them. The proposed class-discriminative attention map (CDAM) is related to both attention and gradient-based explanation methods for deep learning models. On the one hand, explanations are obtained via backpropagation of gradients, but on the other hand, we show that the CDAM scores are directly proportional to the attention.
Gradient-based methods operate by backpropagating gradients from the prediction score to the input features to produce feature attribution maps, also called saliency maps. Many variants have been proposed, see [12, 13, 14, 15, 16] to just name a few. In particular, CDAM is similar to and inspired by the input×$\times$×gradient method [16], because it involves calculating the gradients with respect to tokens multiplied by their activation (Eqs. 2 and 4). The main difference to the input×$\times$×gradient method, and most of other gradient methods, is that we don’t backpropagate the gradients all the way to the input tokens. CDAM stops backpropagation at the tokens that enter the final attention layer. In that regard, our method is related to GradCam [14] which backpropagates the gradients to the final convolutional feature map in a CNN. GradCam and CDAM therefore share the property that they operate on high-level features which is presumably the reason why they are both sensitive to the targeted class, a feature that many explanation methods do not have [17]. Several methods have attempted to apply gradient-based methods or Layer-Wise Relevance Propagation (LRP) [18] to Transformer architectures [19, 20, 21, 22, 23], mostly in the context of NLP tasks.
Attention has often been used to gain insight into a model’s inner workings [6, 7, 8, 9] since it seems intuitive that this weighing of individual tokens correlates with how important they are for the performed task. There is also an ongoing discussion about whether attention provides meaningful explanations [8, 24, 25]. Furthermore, methods such as “attention rollout” [22] have been proposed that aim to quantify how the attention flows through the model to the input tokens. [2] defines the relevancy matrices, as the Hadamard product of the attention map and the gradient of $f_{c}$ w.r.t. to the attention layer, accumulated at each layer. This relevance propagation (RP) method [2], to the best of our knowledge, shows the highest degree of class-discrimination in ViT models. This is used in our study as a comparison for our method.
Many papers have been published on the problem of evaluating post-hoc explanation methods, leading even to the proposal of meta-evaluations [26]. In a recent overview of quantitative evaluation methods [27] 12 desirable properties have been identified. We primarily focus on three of them, namely correctness, contrastivity, and compactness. Our evaluation of correctness is a variant of Sensitivity-$n$ [28], which repeatedly masks $n$ randomly selected pixels and compares the sum of their relevance scores with the change in model output. In the context of ViTs, instead of arbitrarily perturbing observed pixels to some non-informative values, our method called token ablation map (TAM) removes patch tokens before they enter the first transformer block, followed by measuring the resulting difference in model performance. We consider TAM as an intuitive first-order approximation that serves as a sanity check for CDAM. Compactness is measured by looking at the shrinkage and sparsity of CDAM compared to other methods.
3 Proposed Methods
Figure 2: Graphical scheme of class-discriminative attention maps (CDAM). (a) CDAMs are obtained by first propagating an image through the transformer blocks and using its latent representation $l=\texttt{[CLS]}^{\prime}$ to infer a class score $f_{c}$ or similarity score $g(l,l_{c})$. The gradients wrt the layer-normalized input tokens $T_{i}$ of the final attention layer and their activations are then used to calculate the relevance score $S_{i,c}$ according to Eq. 2 or Eq. 4. (b) Detailed view of the final transformer block, adapted from [4].
[…]
In the first scenario, the class score is obtained from the corresponding activation in the prediction vector of a classifier trained on top of the ViT in a supervised fashion. In the second scenario for the concept similarity score, we define a concept through a small set of examples (e.g. ten images of a dog) and obtain a concept vector by averaging their latent representations. Then, without training a classifier, a similarity measure between the concept vector and the latent representation of a targeted sample image is computed and used as the target for the gradients. We call the resulting feature attribution maps “class-discriminative attention maps (CDAM)” because we observe that they scale the attention scores
$A_{i}=\text{softmax}\left(\frac{Q_{\texttt{[CLS]}}K^{\top}_{i}}{\sqrt{d_{k}}}\right),$ (1)
by how relevant the corresponding tokens $T_{i}$ are for the class or similarity score. $Q$ and $K$ are the query and key matrices, consisting of a set of $d_{k}$ dimensional vectors [3]. We discuss the proportionality of CDAM to these attention scores in Section 5.1.
3.1 Relevance Score from Classifiers
[…]
$S_{i,c}=\sum_{j}T_{ij}\frac{\partial f_{c}}{\partial T_{ij}},$ (2)
where $T_{ij}$ is the $j$-th component of the $i$-th token that is fed to the last attention layer [^1] We found in our experiments that calculating the gradients wrt the layer-normalized tokens yields far better results than using the unnormalized ones. (Fig. 2). This definition is a first-order estimation of the contribution of $T_{i}$ to $f_{c}$ in the sense that it would be exact if $f_{c}$ depended on $T_{ij}$ only linearly. It is inspired by gradient-based explanation methods with the difference that we don’t backpropagate the gradients all the way to the input layer but only to the input tokens to the final attention layer. This allows us to gain insight into the decision-making process of the combination of ViT and classifier based on high-level features.
This approach leads to attention maps that are class-discriminative and well-aligned with the semantic relationships of the class and different parts of an image (Fig. 1). As a side note, the heatmaps shown in this article are all scaled independently, thus there is no meaningful comparison of the scale of relevance between images.
3.2 Relevance Score from Similarity Measures
Another way of probing high-level features learned by the ViT is through the use of a similarity measure $g$ in the latent space of the ViT. We compare the latent representation $l(X)=T_{\texttt{[CLS]}^{\prime}}(X)$ of a sample image $X$ and a concept vector $l_{c}$ [29, 10, 11], i.e. the averaged latent representation of $n$ images that are chosen to represent the concept of interest:
$l_{c}=\frac{1}{n}\sum_{X\in c}l(X).$ (3)
This approach is directly inspired by the concept saliency map [10]. The definition of the relevance score in this setting is analogous to Eq. 2
$S_{i,c}=\sum_{j}T_{ij}\frac{\partial g(l,l_{c})}{\partial T_{ij}}.$ (4)
Possible choices for $g$ are for example the $L^{2}$ distance, cosine similarity or dot product, where we use the latter in our experiments. The definition Eq. 4 estimates the contribution of $T_{i}$ to $g$, i.e. it allows us to gauge how much each token is driving the (dis-)similarity between $l$ and $l_{c}$.
3.3 Token Ablation Study
As an independent comparison method, we carried out a token ablation study, in which a change in the model accuracy is measured after removing one token at a time. We visualize changes in the model accuracy as token ablation maps (TAM). This method for ViT is analogous to input pixel occlusion [30]. Such occlusion methods gauge a (group of) pixel’s relevance by masking them with non-informative values and recording the change in model output. This approach is problematic because perturbing pixels potentially leads to out-of-distribution samples which may trigger artifacts of the model [31, 32]. Relatedly, in contrast to pixel occlusion methods, for TAM there is no need to choose a specific way of perturbing input features in pixel space, which would lead to some ambiguity and variation in the resulting relevance maps. The token ablation study sidesteps this issue in the context of ViTs by removing image patch tokens, one at a time, before they enter the first transformer block, instead of masking image patches in pixel space. The relevance score assigned by TAM is defined by
$S^{\text{TAM}}_{i,c}=f_{c}(\mathbf{T})-f_{c}(\mathbf{T}\backslash{T_{i}}),$ (5)
where $\mathbf{T}$ is the set of all tokens that enter the first transformer block and $\mathbf{T}\backslash{T_{i}}$ has token $T_{i}$ removed from $\mathbf{T}$.
Note that TAMs are not to be confused with the “ground truth” of feature relevance since we are dealing with non-linear models with complex interactions of patch tokens. Therefore, tokens would likely be assigned different relevance scores if we perturbed two or more tokens simultaneously, with different scores depending on the chosen group of tokens. This effect has previously been demonstrated in the context of CNNs [28]. Therefore, comparing with TAM mainly serves as a sanity check for CDAM, since TAM is a first approximation to patch token relevance with a clearly defined meaning.
4 Results
Figure 3: Class-Discriminative Attention Maps (CDAM) for user-defined concepts (Section 3.2). The concept embedding $l_{c}$ has been obtained by averaging the latent representations of 30 images that include the common concept. In each instance, from left to right: Sample image, AM, and CDAM. Orange and blue correspond to positive and negative values, respectively.
For our experiments, we have been using the ViT-S/8 architecture trained with DINO as a backbone [1], which is available as a pre-trained model in their repository [^2] https://github.com/facebookresearch/dino.
4.1 Relevance Scores from Classifier
To demonstrate our method, we have trained a linear classifier on the ImageNet [33] that achieved an accuracy of $77.0\%$. We use random resized cropping and horizontal flipping with PyTorch default arguments as augmentation, Adam optimizer with learning rate $3\times 10^{-4}$, batch size 128, and train for 10 epochs. The parameters of the ViT backbone are frozen during training, so only the classifier head is trainable. The predictions of this classifier were then used to create the heat maps shown in Fig. 1 via Eq. 2. Notice how they clearly distinguish the targeted class with positive values (evidence) from the rest of the image, which is assigned either zero (irrelevant) or negative scores (counter-evidence). Note that CDAM does not assign non-zero $S_{i,c}$ to tokens that have zero attention in AM. CDAM therefore appears to compute $S_{i,c}$ by multiplying the attention with the relevance of the corresponding token for the targeted class. This is a nice feature since CDAM inherits the high-quality object segmentations present in AM, which are focused on the objects of interest. We discuss this operating characteristic in Section 5.
One of the most interesting features of CDAM is that it is highly sensitive to the chosen class (Fig. 4, left). For example, when either the zuccini or bell pepper class is chosen, CDAM clearly highlights that vegetable. Most of the other vegetables are assigned either near-zero or negative values. Interestingly, the green stems of other vegetables are also assigned positive values. CDAM reveals that the ViT has learned to associate input features with chosen classes that are largely consistent with human expectations. Additional examples of CDAM and other interpretability methods are showned in the Appendix A (Figs. 7, 8, 9 and 10).
Figure 4: When targeting different classes (left) or concepts (right), the resulting CDAMs are distinct and align with the corresponding objects. Note that green parts of vegetables are shown to activate the zuccini class (left). The wrinkles in the elephant’s trunk appear to be related to the zebra concept, probably semantically mistaken by the model as zebra stripes (zoom in for a better view).
4.2 Relevance Scores from Similarity Measure
Concept embeddings $l_{c}$ have been obtained by averaging 30 latent representations of images depicting the concept of interest. These images were randomly selected from the corresponding classes in the ImageNet validation set [33] and the CDAM in Fig. 3 were obtained using Eq. 4. Again we observe that CDAM reveals that the model clearly distinguishes the parts of the images that are relevant for the chosen concept from the rest of the image. For instance, for the concept hammer, CDAM highlights not only the hammer itself but also the wooden handles of the screwdrivers and the top of the ratchet, which are indeed relevant for this concept. Unrelated tools or parts appear in blue indicating that the model deems them to be not similar to the concept hammer.
Just as CDAM is sensitive to the chosen class, it also produces attention maps that are specific to the chosen concept (Fig. 4, right). This makes the proposed method particulary valuable for understanding the representations the ViT has learned since the explanations for different concepts can give a complementary view. For instance, CDAM shows that parts of the elephant’s trunk are counter-evidence for the concept african elephant (Fig. 4, right) which is at first counter-intuitive. However, when compared with the concept zebra, we can see that the model deems this part of the trunk to be similar to a zebra and it probably mistakes the trunk’s wrinkles for the zebra’s stripes. By choosing a set of images with a common concept, CDAM helps probing the latent space of the ViT, allowing us to learn about how it “sees” the world in a semi-supervised manner.
4.3 Comparison with relevance propagation and token ablation maps
Figure 5: Comparison of specificity to the target class. AM is, by definition, not influenced by it, but relevance propagation [2] does show some specificity. In particular due to the signed relevance scores both CDAM and TAM exhibit strong class-sensitivity and the two maps roughly agree, with CDAM appearing as stronger regularized.
Upon visual inspection, it appears that CDAM is estimating patch token relevance that aligns well with our expectation concerning the most relevant (counter-)evidence for a given target class (Figs. 1 and 4). To investigate whether these relevance scores correspond to the actual inference performed by the model, we compare them to token ablation maps (TAMs) and relevance propagation (RP) [2]. In TAMs, the model output is measured after an individual token is removed (Section 3.3). RP, which is based on the Deep Taylor Decomposition, is an explanation method for ViTs that, to the best of our knowledge, produces relevance scores that are the most class-discriminative so far [2, 23].
We find that RP is more class-discriminative than AM and that CDAM and TAM distinguish the different classes more clearly than RP ( Fig. 5). Visually, the signed scores in CDAM and TAM make stronger distinctions between the different classes, whereas RP only “dims” the object that is not targeted. Fig. 6 shows pair-wise comparisons among scores for CDAM, TAM, AM, and RP. We observe that RP tends to assign low but positive scores to tokens that obtained negative scores from CDAM or TAM.
[…]
Compared to AM, TAM, and RP, CDAM appears to be less noisy (Fig. 5). We can observe this implicit regularization effect by looking at the scores from 1000 randomly selected images Fig. 6. In AM, TAM, and RP, the per-image average percentage of patch tokens with relevance scores smaller than $5\%$ of the maximum score are $54\%,66\%,49\%$, respectively, whereas in CDAM it is $84\%$. While sparsity may not always imply the most accurate estimation at a single data point (e.g., a pixel), regularization has long been pursued and motivated by interpretability in machine learning [34].
Figure 6: Pairwise scatter plots of relevance/attention scores for CDAM, TAM, AM, and relevance propagation obtained from 1000 randomly selected images from the ImageNet validation set. The target class for the explanations is the one with the highest logit score. Notice in particular that AM and relevance propagation are highly correlated and that CDAM and TAM largely agree on the sign of relevance scores. We also observe that $A_{i}=0\Rightarrow S_{i,c}=0$, where $A_{i}$ and $S_{i,c}$ refer to the AM or CDAM score of token $T_{i}$.
5 Discussion
Class-Discriminative Attention Map (CDAM) is a gradient-based extension of attention maps (AM), that provides strong discrimination with respect to a chosen class or concept. In contrast, many post-hoc explanation methods produce similar heatmaps even when targeting different classes [23, 17]. CDAM retains appealing characteristics of AM, namely high-quality semantic segmentation in self-supervised ViTs [1]. Previously, a considerable effort has been made to de-noise post-hoc interpretability methods by averaging variations of multiple saliency maps [13, 15]. Without further processing or averaging, the proposed CDAM produces comparatively sparse heat maps that are less noisy than previous methods.
Besides explaining predictions of the classifier on the top of ViT, we also introduce how to obtain a CDAM by measuring the similarity between the latent representation of a sample image and a concept embedding. By probing the latent space of the ViT with different concepts (Fig. 4) we can learn a lot about the representations it has learned. To some degree this can of course also be done by looking at the attention maps of the different heads, which tend to specialize in different categories of objects or parts thereof [1]. This approach is, however, limited to the categories that the attention heads happened to have learned. CDAM allows us to probe for concepts that are relevant to any given use case.
5.1 Proportionality of CDAM to the attention
CDAM appears to scale the attention $A_{i}=\text{softmax}\left(\frac{Q_{\texttt{[CLS]}}K^{\top}_{i}}{\sqrt{d_{k}}}\right)$ by the relevance of the corresponding token $T_{i}$ for a concept. To better understand this observation, without loss of generality, consider a single attention head. Eq. 4 is expanded
$S_{i,c}$ $=\sum_{j}T_{ij}\frac{\partial g(l,l_{c})}{\partial T_{ij}}$
$=\sum_{j}T_{ij}\frac{\partial}{\partial T_{ij}}\sum_{k}l_{k}l_{c,k}$
$=\sum_{j}T_{ij}\frac{\partial}{\partial T_{ij}}\sum_{k}h_{k}\left(\sum_{m}A_{m}V_{m}+\texttt{[CLS]}\right)l_{c,k}$
$=\sum_{j}T_{ij}\sum_{k,m,n}\nabla_{n}h_{k}\times\left(\frac{\partial A_{m}}{\partial T_{ij}}V_{mn}+A_{m}\frac{\partial V_{mn}}{\partial T_{ij}}\right)l_{c,k}$
$=\sum_{j,k,m,n}T_{ij}\nabla_{n}h_{k}\frac{\partial A_{m}}{\partial T_{ij}}V_{mn}l_{c,k}+\sum_{k,n}A_{i}V_{in}\nabla_{n}h_{k}l_{c,k},$ (6)
[…]
Importantly, we observe in our multi-head attention ViT model (Fig. 6) that
$A_{i}=0\Rightarrow S_{i,c}=0$ (7)
which implies that in Section 5.1, all terms that are not proportional to $A_{i}$ are zero or cancelled out and we are left with
$S_{i,c}=A_{i}V_{i}\cdot\nabla(h\cdot l_{c}).$ (8)
Empirically, $S_{i,c}$ is obtained by multiplying $A_{i}$ with the directional derivative of $h\cdot l_{c}$ in the direction of $V_{i}$. In other words, CDAM scales $A_{i}$ by the rate at which the dot product (similarity) between $h$ and $c$ changes in the direction of $V_{i}$. In the case that $h$ is just the identity function, we obtain an even simpler interpretation of CDAM. That’s because we then have $\nabla_{n}h_{k}(f)=\partial_{f_{n}}f_{k}=\delta_{nk}$ and get from Section 5.1, again assuming $\frac{\partial A_{m}}{\partial T_{ij}}=0$, that $S_{i,c}=A_{i}V_{i}\cdot l_{c}$, which means that $A_{i}$ is scaled by the similarity of $V_{i}$ and $l_{c}$.
5.2 Faithfulness of Post-hoc Explanation Methods
Understanding the decision-making process of deep neural networks is intrinsically challenging due to the vast parameter space and non-linearity, which lead to complex interactions among the patch tokens in ViT. Therefore, post-hoc explanation methods such as RP, CDAM, and TAM can only explain some facets of the model and can never be comprehensively faithful to it, as that would entail being equivalent to the model itself [17]. Of course, the complete set of the model weights in DNNs would not be considered interpretable. Therefore, attempting to create explanation methods necessitates a balance among many properties, ranging from sparsity to discriminativeness [27].
The proposed CDAM relies on gradients and is therefore a first-order approximation. Instead of referring directly to the input patch tokens, the CDAM scores refer to tokens that, in general, have been enriched with contributions from all other tokens through the attention layers. Relevance is therefore not directly attributed to input features. Nonetheless, by extending the inherent attention scores in ViTs, CDAM captures the high-quality semantic segmentations of input images shown in self-supervised ViTs [1], while providing class-discrimination.
6 Conclusion
By extending the attention maps (AM) inherent to ViT, the proposed CDAM provides relevance scores for a chosen class or concept. The proposed relevance estimation for CDAM essentially scales the attention scores by how relevant they are for a target class or concept. As a result, zero attention scores remain zero in CDAM. The resulting heat maps clearly distinguish the object representing the target from the background, as well as other objects that are present. In addition to using a classifier, we proposed to explain a user-defined concept by selecting a small set of example images. Using similarity measures in the latent representation, CDAM provides highly specific heat maps for a concept that underly selected examples. Especially in the context of SSL where class labels are absent, this makes it a valuable technique to investigate the model representations.
In terms of class-discrimination, our method compares favorably against the state of the art, namely relevance propagation (RP) for transformers [2]. Especially, the positive and negative signs of relevance scores in CDAM make clear distinctions qualitatively and quantitatively. We also devise the token ablation maps (TAM) which visualize the changes in the ViT model output when corresponding image patch tokens are removed. The advantage of TAM is its explicable computation, which serves as a sanity check. CDAM captures the overall patterns of TAM while providing sparser and more semantically consistent heat maps. However, while RP and TAM are used in comparison, neither of them is the “ground truth” of relevance scores.
Post-hoc explanation methods for DNNs approximate the full operating characteristics of the model with certain desirable properties. Our results suggest that CDAM is capable of achieving a high degree of correctness, contrastivity, and compactness, as described in [35]. Note that CDAM is not explicitly designed to satisfy these criteria. For example, we could impose further sparsity constraints that would increase compactness. However, over-optimizing an explanation method for certain properties could have undesirable effects [^3] Goodhart’s Law states that “when a measure becomes a target, it ceases to be a good measure” [36]. While this paper uses a ViT model trained with SSL [1], CDAM is applicable more broadly to all transformer models with self-attention, e.g., large language models [37].
Acknowledgments
This work was funded by the ERA-Net CHIST-ERA grant [CHIST-ERA-19-XAI-007] long term challenges in ICT project INFORM (ID: 93603), by the National Science Centre (NCN) of Poland [2020/02/Y/ST6/00071]. This research was carried out with the support of the Interdisciplinary Centre for Mathematical and Computational Modelling University of Warsaw (ICM UW) under computational allocation no GDM-3540; the NVIDIA Corporation’s GPU grant; and the Google Cloud Research Innovators program.
Appendix A Appendix
A.1 Multiple attention heads
We note that introducing $h$ attention heads does not change the derivations in Section 5.1 qualitatively. Instead of the sum $Z=\sum_{m}A_{m}V_{m}$ there will be a term $\text{Concat}(Z^{1},Z^{2},..,Z^{h})W^{O},W^{O}\in\mathbb{R}^{hd_{v}\times d_{\text{model}}}$ so that in the case of multi-head attention the equivalent of Section 5.1 reads
$S_{i,c}=\text{Concat}\left(Y^{1},Y^{2},..,Y^{h}\right)W^{O}\nabla(h\cdot l_{c}),\;Y^{n}=T_{i}\sum_{m}\frac{\partial A^{n}_{m}}{\partial T_{i}}V^{n}_{m}+A^{n}_{i}V^{n}_{i}$ (9)
and since $A^{n}_{i}>0,A_{i}=\sum_{n}\frac{A_{i}^{n}}{h}$ implies that $A^{i}=0\Rightarrow A_{i}^{n}=0,\forall n$ we can again conclude that in order for Eq. 7 to hold all terms that are not proportional to $A_{i}$ either vanish or cancel out. For multi-head attention we thus arrive at
$S_{i,c}=\text{Concat}\left(Y^{\prime 1},Y^{\prime 2},..,Y^{\prime h}\right)W^{O}\nabla(h\cdot l_{c}),\;Y^{\prime n}=A^{n}_{i}V^{n}_{i},$ (10)
which does not have such a straightforward interpretation as the single-headed case due to the mixing of the attention heads.
If we think of $W^{O}$ as acting first on $\nabla(h\cdot l_{c})$, creating a column vector of dimension $hd_{v}$, we can understand Eq. 10 as the sum of $h$ directional derivatives of the $d_{v}$-long segments of this columns vector in the directions of $Y^{\prime 1}$ to $Y^{\prime h}$. Each of the summands is proportional to the corresponding $A^{n}$, and the interpretation of CDAM from the single-head case therefore extends to the multi-head one.
Although our observation $A_{i}=0\Rightarrow S_{i,c}=0$ implies that all terms that are not proportional to $A_{i}$ in Section 5.1 should vanish, we don’t have a theoretical argument why that would be the case. It seems unlikely that the various terms in the first sum cancel out, which leaves $\frac{\partial A_{m}}{\partial T_{ij}}=0$ as an explanation. However, in general, $A_{i}$ depends on $T_{i}$ through $K_{i}$ and one would therefore expect the derivative to be non-zero.
We can also show that CDAM scales the attention by the relevance of the correpsonding token for a chosen class (instead of concept), in the case that a classifier with weights $W^{C}$ has been trained on top of the ViT. Analogously to the above single-head discussion we have
$S_{i,c}$ $=\sum_{j}T_{ij}\frac{\partial f_{c}}{\partial T_{ij}}$
$=\sum_{j}T_{ij}\frac{\partial}{\partial T_{ij}}\sum_{k}l_{k}W^{C}_{kc}$
$=\sum_{j,k,m,n}T_{ij}\nabla_{n}h_{k}\frac{\partial A_{m}}{\partial T_{ij}}V_{mn}W^{C}_{kc}+\sum_{k,n}A_{i}V_{in}\nabla_{n}h_{k}W^{C}_{kc},$ (11)
and the only difference is that the directional derivative now acts on $\sum_{k}h_{k}W_{kc}$, i.e. the contribution of $h$ to the activation of class $c$ in the prediction vector, instead of $\sum_{k}h_{k}l_{c,k}$. All the intuitions we have gained from the concept-based CDAM therefore extend to the class-based one.
A.2 Examples for explanation methods
Figure 7: Examples for the considered explanation methods. The indicated class is the one predicted by the model and is the target for the explanation methods.
Figure 8: Examples for the considered explanation methods. The indicated class is the one predicted by the model and is the target for the explanation methods.
Figure 9: Examples for the considered explanation methods. The indicated class is the one predicted by the model and is the target for the explanation methods.
Figure 10: Examples for the considered explanation methods. The indicated class is the one predicted by the model and is the target for the explanation methods.

### v2

@@
 Abstract
-Interpretability methods are critical components for examining and exploring deep neural networks (DNN), as well as increasing our understanding of and trust in them.
-Vision transformers (ViT), which can be trained to state-of-the-art performance with a self-supervised learning (SSL) training method [1], provide built-in attention maps (AM).
-While AMs can provide high-quality semantic segmentation of input images, they do not account for any signal coming from a downstream classifier.
-We introduce class-discriminative attention maps (CDAM), a novel post-hoc explanation method that is highly sensitive to the target class.
-Our method essentially scales attention scores by how relevant the corresponding tokens are for the predictions of a classifier head.
-Alternative to classifier outputs, CDAM can also explain a user-defined concept by targeting similarity measures in the latent space of the ViT.
-This allows for explanations of arbitrary concepts, defined by the user through a few sample images.
-We investigate the operating characteristics of CDAM in comparison with relevance propagation (RP) [2] and token ablation maps (TAM), an alternative to pixel occlusion methods.
-CDAM is highly class-discriminative and semantically relevant, while providing implicit regularization of relevance scores.
+Importance estimators are explainability methods that quantify feature importance for deep neural networks (DNN).
+In vision transformers (ViT), the self-attention mechanism naturally leads to attention maps, which are sometimes used as importance scores for which input features ViT models are focusing on.
+However, attention maps do not account for signals from downstream tasks.
+To generate explanations that are sensitive to downstream tasks, we have developed class-discriminative attention maps (CDAM), a gradient-based extension that estimates feature importance with respect to a known class or a latent concept.
+CDAM scales attention scores by how relevant the corresponding tokens are for the predictions of a classifier head.
+In addition to targeting the supervised classifier, CDAM can explain an arbitrary concept shared by selected samples by measuring similarity in the latent space of ViT.
+Additionally, we introduce Smooth CDAM and Integrated CDAM, which average a series of CDAMs with slightly altered tokens.
+Our quantitative benchmarks include correctness, compactness, and class sensitivity, in comparison to six other importance estimators.
+Vanilla, Smooth, and Integrated CDAM excel across all three benchmarks.
+In particular, our results suggest that existing importance estimators may not provide sufficient class-sensitivity.
+We demonstrate the utility of CDAM in medical images by training and explaining malignancy and biomarker prediction models based on lung Computed Tomography (CT) scans.
+Overall, CDAM is shown to be highly class-discriminative and semantically relevant, while providing compact explanations.
 PyTorch implementation: https://github.com/lenbrocki/CDAM.
 Web live demo: https://cdam.informatism.com/.
-Figure 1: Extending attention maps (AM), the proposed class-discriminative attention maps (CDAM) quantify and visualize input features that are relevant for the target class in Vision Transformers (ViT).
-We trained a linear classifier on the ViT-S/8 trained with DINO as a backbone [1] and visualize random samples with CDAM for different target classes.
-For details, refer to Section 3.1.
-Orange and blue colors correspond to positive and negative values, respectively.
 1 Introduction
-The transformer architecture, which was originally developed for natural language processing [3], has been developed into vision transformers (ViT) that demonstrate comparable or even better performance than convolutional neural networks (CNN) on many computer vision tasks [4, 5].
-Additionally, a powerful feature of transformers is that the attention mechanism provides a built-in explanation of the model’s inner workings, if attention is understood as a weighing of the relevance of individual tokens [6, 7, 8, 9].
-While attention maps (AM) provides intuitive interpretability, it fails to be discriminative with respect to a target class.
-Recognizing the potential of AM for explaining a ViT’s inner workings, we propose class-discriminative attention maps (CDAM), a gradient-based extension that makes them sensitive to class-specific signals coming from a downstream classifier.
-Recently, it has been demonstrated that ViT models trained with DINO, a self-supervised learning (SSL) training method [1], produce AMs that provide high-quality semantic segmentations of input images.
-These AMs clearly distinguish objects of interest from the background with very little noise, with different attention heads often focusing on semantically distinct regions [1].
-AMs are the visualization of the self-attention of the [CLS] token query in the final attention layer.
-However, one shortcoming of AMs is that they purely operate on the level of the last attention layer and therefore do not take into account any information coming from a downstream task, such as a classifier head.
-The proposed CDAMs rely on the calculation of gradients of the class score, i.e.
-the logit value of a target class, with respect to the token activations in the final transformer block (Fig.
+Vision transformers (ViT) are a family of computer vision models based on the transformer architecture [1], that demonstrate comparable or even better performance than convolutional neural networks (CNN) on many object detection and classification tasks [2, 3].
+A powerful feature of the transformers is that the attention mechanism provides a built-in explanation of the model’s inner workings, if attention is understood as a weighing of the relevance of individual tokens [4, 5, 6, 7].
+While attention maps provide intuitive and inherent interpretability, they fail to be discriminative with respect to a target class.
+Recognizing the potential of attention maps for explaining a ViT’s inner workings, we have developed class-discriminative attention maps (CDAM), a gradient-based extension that makes them sensitive to class-specific signals based on a classifier or a concept derived from selected samples.
+Attention maps are the visualization of the self-attention of the class token query [CLS] in the final attention layer.
+Attention heads can be visualized separately, and in certain instances, may be related to semantically meaningful concepts (e.g., in language [8]).
+Recently, ViT trained with self-supervised learning (SSL) methods such as DINO [9] have shown excellent performance leveraging a greater amount of unlabeled data.
+Attention maps from such self-supervised ViT models provide high-quality semantic segmentations of input images.
+These attention maps clearly distinguish regions of interest (ROI) from the background with very little noise, with different attention heads often focusing on semantically distinct regions [9].
+There have been many improvements to ViT including the introduction of convolution [10, 11] and spatial sparsity [12].
+Attention maps, as well as our proposed CDAMs, are applicable to all ViT architectures with self-attention.
+The major shortcoming of attention maps, when considering them as an explainable AI (XAI) method, is that they purely operate on the level of the last attention layer and therefore do not take into account any information coming from a downstream task, such as a classifier head.
+The proposed CDAM overcomes this shortcoming by calculating the gradients of the class score, i.e.
+the logit value of a target class, with respect to the token activations in the final transformer block (visualized in Fig.
 2(a)).
-We show through explicit calculation and empirical observations that CDAM scales the attention by how relevant the corresponding tokens are for the model’s decision.
-CDAM therefore inherits the desirable properties of AMs, but additionally indicates the (counter-)evidence for a target class that the ViT relies on for its predictions (Fig.
+Essentially, CDAM scales the attention scores by how relevant the corresponding tokens are for the model’s decision.
+CDAM therefore inherits the desirable properties of attention maps, but additionally indicates the evidence or counter-evidence for a target class that the ViT relies on for its predictions (examples in Fig.
 1).
 We demonstrate that it is class-discriminative in the sense that the heat maps clearly distinguish the object representing the target class from the background and, importantly, from objects that belong to other classes.
-CDAMs are also highly sensitive to the target class which, all together, makes it a useful tool for explaining the decision-making process of classifiers with a ViT backbone.
-In addition to the classifier-based explanations, we present a concept-based approach that replaces the class score with a similarity measure in the latent space of the ViT, where a concept is defined through a number of example images.
+Therefore, CDAM is a useful tool for explaining the decision-making process of classifiers with a ViT backbone.
+In addition to the classifier-based explanations, we present a concept-based approach that replaces the class score with the similarity measure in the latent space of the ViT, where a concept is defined through selected images.
 Averaging the latent representations of the example images yields a concept embedding, which is analogous to a concept vector in variational autoencoders (VAE) and related generative models [13, 14].
-In brief, we use CDAM to identify image patches that drive the (dis-)similarity of the concept embedding and the latent representation of a given image.
 This allows obtaining a heat map for the target concept the model has never been explicitly trained on, which is of particular interest for self-supervised models.
-Both supervised and semi-supervised variants of CDAM are compared to token ablation maps (TAMs) and relevance propagation (RP) [2].
-Overall, our results suggest that CDAMs exhibit semantic coherency, implicit regularization, and strong discrimination with respect to a target class/concept.
+We further demonstrate when computing gradients of a class (Section 3.1) or concept similarity (Section 3.2) score, with respect to the token activations in the final transformer block, smoothing or averaging could be applied to improve performance.
+Specifically, we introduce two additional variants of CDAM; (1) Smooth CDAM averages over multiple gradients that are computed with additional noise to tokens (inspired by SmoothGrad [15]) and (2) Integrated CDAM takes the intergral of the gradients along the path from the baseline to the token in the final transformer block (inspired by IntGrad [16]).
+We conduct several quantitative evaluations focusing on correctness, class sensitivity, and compactness.
+Correctness is measured by perturbing an increasing amount of input features and measuring the model output [17, 18].
+By using the ImageNet samples [19] with multiple objects [20] and applying importance estimators for different classes, we quantify the level of class-discrimination.
+Sparsity and shrinkage of explanations are measured by the number of near-zero importance scores.
+Besides vanilla, Smooth, and Integrated CDAM, we also employed attention maps, relevance propagation (RP) [21], partial Layer-wise Relevance Propagation (LRP) [22], Input $\times$ Gradients, SmoothGrad [15], and Integrated Gradient [16] in comparison.
+Three variants of CDAM outperform other explainability methods on three benchmarks.
+Lastly, we have applied CDAM on a ViT fine-tuned on the Lung Image Database Consortium image collection (LIDC) [23].
+Two ViT-based models are built for predicting malignancy and biomarkers.
+CDAM, as well as attention maps, are obtained.
+CDAM is shown to be highly class-discriminative and semantically relevant, while providing implicit regularization of importance scores.
 2 Related Works
-Interpretability of ViT, and more broadly deep neural networks (DNN), is an active area of research, since understanding their decision-making process would help improve the model, identify weaknesses and biases, and ultimately increase human trust in them.
+Interpretability of ViT, and more broadly explainable artificial intelligence (XAI), is an active area of research, since understanding their decision-making process would help improve the model, identify weaknesses and biases, and ultimately increase human trust in them.
 The proposed class-discriminative attention map (CDAM) is related to both attention and gradient-based explanation methods for deep learning models.
-On the one hand, explanations are obtained via backpropagation of gradients, but on the other hand, we show that the CDAM scores are directly proportional to the attention.
-Gradient-based methods operate by backpropagating gradients from the prediction score to the input features to produce feature attribution maps, also called saliency maps.
+Gradient-based methods operate by backpropagating gradients from the prediction score to the input features to produce feature attribution maps, which is also called saliency maps.
 Many variants have been proposed, see [24, 16, 25, 15, 26] to just name a few.
@@
 2 and 4).
-The main difference to the input×$\times$×gradient method, and most of other gradient methods, is that we don’t backpropagate the gradients all the way to the input tokens.
+The main difference to most gradient-based methods is that we don’t backpropagate the gradients all the way to the input tokens.
 CDAM stops backpropagation at the tokens that enter the final attention layer.
@@
 This relevance propagation (RP) method [21], to the best of our knowledge, shows the highest degree of class-discrimination in ViT models.
-This is used in our study as a comparison for our method.
+This is used in our study as one of comparative importance estimators.
 Many papers have been published on the problem of evaluating post-hoc explanation methods, leading even to the proposal of meta-evaluations [35].
@@
 We primarily focus on three of them, namely correctness, contrastivity, and compactness.
-Our evaluation of correctness is a variant of Sensitivity-$n$ [28], which repeatedly masks $n$ randomly selected pixels and compares the sum of their relevance scores with the change in model output.
-In the context of ViTs, instead of arbitrarily perturbing observed pixels to some non-informative values, our method called token ablation map (TAM) removes patch tokens before they enter the first transformer block, followed by measuring the resulting difference in model performance.
-We consider TAM as an intuitive first-order approximation that serves as a sanity check for CDAM.
-Compactness is measured by looking at the shrinkage and sparsity of CDAM compared to other methods.
-3 Proposed Methods
-Figure 2: Graphical scheme of class-discriminative attention maps (CDAM).
-(a) CDAMs are obtained by first propagating an image through the transformer blocks and using its latent representation $l=\texttt{[CLS]}^{\prime}$ to infer a class score $f_{c}$ or similarity score $g(l,l_{c})$.
-The gradients wrt the layer-normalized input tokens $T_{i}$ of the final attention layer and their activations are then used to calculate the relevance score $S_{i,c}$ according to Eq.
-2 or Eq.
-4.
-(b) Detailed view of the final transformer block, adapted from [4].
+In particular, carefully designed occulusion/perturbation studies can help us calculate a fidelity statistics [17], which is also called the symmetric relevance gain (SRG) measure [37].
+[38] first demonstrated that many importance estimators, which purport to be sensitive to the internals of a classifier, produce similar importance scores regardless of the relation of inputs and classes in the training data, which they called “the data randomization test”.
+Generally, if two importance scores with respect to two different classes produce very similar heat maps, the importance estimator is not class sensitive.
+We quantify this concept in our class sensitivity evaluation benchmark to measure whether an explanation is sensitive to the classes.
+3 Methods
+Adaptation of the transformer architecture to computer vision hinges on creating tokens from small patches of the input image [2].
+In addition to those image patch tokens, the [CLS] token is added, which is utilized for downstream tasks such as classification.
+Attention maps are the two-dimensional visualization of the self-attention in the final attention layer with [CLS] as query token, where the attention scores are averaged over all heads.
 […]
 In the first scenario, the class score is obtained from the corresponding activation in the prediction vector of a classifier trained on top of the ViT in a supervised fashion.
-In the second scenario for the concept similarity score, we define a concept through a small set of examples (e.g.
+In the second scenario, we define a concept through a small set of examples (e.g.
 ten images of a dog) and obtain a concept vector by averaging their latent representations.
 Then, without training a classifier, a similarity measure between the concept vector and the latent representation of a targeted sample image is computed and used as the target for the gradients.
-We call the resulting feature attribution maps “class-discriminative attention maps (CDAM)” because we observe that they scale the attention scores
+We call the resulting maps of feature relevance “class-discriminative attention maps” because they scale the attention scores
 $A_{i}=\text{softmax}\left(\frac{Q_{\texttt{[CLS]}}K^{\top}_{i}}{\sqrt{d_{k}}}\right),$ (1)
@@
 $Q$ and $K$ are the query and key matrices, consisting of a set of $d_{k}$ dimensional vectors [1].
-We discuss the proportionality of CDAM to these attention scores in Section 5.1.
-3.1 Relevance Score from Classifiers
+We discuss the proportionality of CDAM to the attention scores in Section A.1.
+Self-supervised learning (SSL) methods such as DINO can train ViT models whose attention maps are high-quality semantic segmentations of the input, without using labels [9].
+Therefore, we utilize a ViT model [Dosovitskiy2020ViT] pre-trained with DINO as a backbone for our experiments and applications.
+We denote details about the ViT architecture, training schemes, and model performance in the Section 4.
+Note that the heat maps shown in this article are scaled independently.
+3.1 CDAM for a known class
 […]
 $S_{i,c}=\sum_{j}T_{ij}\frac{\partial f_{c}}{\partial T_{ij}},$ (2)
-where $T_{ij}$ is the $j$-th component of the $i$-th token that is fed to the last attention layer [^1] We found in our experiments that calculating the gradients wrt the layer-normalized tokens yields far better results than using the unnormalized ones.
-(Fig.
+where $T_{ij}$ is the $j$-th component of the $i$-th token that is fed to the last attention layer (Fig.
 2).
+We found in our experiments that calculating the gradients with respect to the layer-normalized tokens yields far better results than using the unnormalized ones.
 This definition is a first-order estimation of the contribution of $T_{i}$ to $f_{c}$ in the sense that it would be exact if $f_{c}$ depended on $T_{ij}$ only linearly.
@@
 1).
-As a side note, the heatmaps shown in this article are all scaled independently, thus there is no meaningful comparison of the scale of relevance between images.
-3.2 Relevance Score from Similarity Measures
+3.2 CDAM for a latent concept
 Another way of probing high-level features learned by the ViT is through the use of a similarity measure $g$ in the latent space of the ViT.
@@
 $l_{c}=\frac{1}{n}\sum_{X\in c}l(X).$ (3)
-This approach is directly inspired by the concept saliency map [10]. The definition of the relevance score in this setting is analogous to Eq. 2
+This approach is directly inspired by the concept saliency map [13]. The definition of the importance score in this setting is analogous to Eq. 2
 $S_{i,c}=\sum_{j}T_{ij}\frac{\partial g(l,l_{c})}{\partial T_{ij}}.$ (4)
-Possible choices for $g$ are for example the $L^{2}$ distance, cosine similarity or dot product, where we use the latter in our experiments.
+Possible choices for $g$ are for example the $L^{2}$ distance, cosine similarity or dot product, where we use the latter in this article.
 The definition Eq.
-4 estimates the contribution of $T_{i}$ to $g$, i.e.
-it allows us to gauge how much each token is driving the (dis-)similarity between $l$ and $l_{c}$.
-3.3 Token Ablation Study
-As an independent comparison method, we carried out a token ablation study, in which a change in the model accuracy is measured after removing one token at a time.
-We visualize changes in the model accuracy as token ablation maps (TAM).
-This method for ViT is analogous to input pixel occlusion [30].
-Such occlusion methods gauge a (group of) pixel’s relevance by masking them with non-informative values and recording the change in model output.
-This approach is problematic because perturbing pixels potentially leads to out-of-distribution samples which may trigger artifacts of the model [31, 32].
-Relatedly, in contrast to pixel occlusion methods, for TAM there is no need to choose a specific way of perturbing input features in pixel space, which would lead to some ambiguity and variation in the resulting relevance maps.
-The token ablation study sidesteps this issue in the context of ViTs by removing image patch tokens, one at a time, before they enter the first transformer block, instead of masking image patches in pixel space.
-The relevance score assigned by TAM is defined by
-$S^{\text{TAM}}_{i,c}=f_{c}(\mathbf{T})-f_{c}(\mathbf{T}\backslash{T_{i}}),$ (5)
-where $\mathbf{T}$ is the set of all tokens that enter the first transformer block and $\mathbf{T}\backslash{T_{i}}$ has token $T_{i}$ removed from $\mathbf{T}$.
-Note that TAMs are not to be confused with the “ground truth” of feature relevance since we are dealing with non-linear models with complex interactions of patch tokens.
-Therefore, tokens would likely be assigned different relevance scores if we perturbed two or more tokens simultaneously, with different scores depending on the chosen group of tokens.
-This effect has previously been demonstrated in the context of CNNs [28].
-Therefore, comparing with TAM mainly serves as a sanity check for CDAM, since TAM is a first approximation to patch token relevance with a clearly defined meaning.
+4 estimates the contribution of $T_{i}$ to $g$, such that it allows us to gauge how much each token is driving the similarity between $l$ and $l_{c}$.
+3.3 Smooth and Integrated CDAM
+SmoothGrad (SG) [15] and Integrated Gradients (IG) [16] extended saliency maps [24], which were introduced in the context of CNNs as visualizing gradients of the class with respect to the input features.
+Essentially, these two modifications sharpen and denoise vanilla saliency maps, with theoretical and practical justifications.
+Briefly, [15] adds a small Gaussian noise to a sample image $n$ times, resulting in $n$ vanilla saliency maps.
+SG is an average of those $n$ saliency maps.
+In IG, instead of adding random noise, [16] uses multiple saliency maps obtained from interpolating between a given sample and a baseline (e.g., a black image).
+We have adapted the analogous modifications to CDAM, resulting in two additional variants which we call Smooth and Integrated CDAM.
+Instead of modifying the input images, we vary the tokens that enter the final transformer block.
+In Integrated CDAM, we interpolate between a baseline and the tokens themselves, resulting in $n$ maps.
+In Smooth CDAM, we repeatedly add a small noise to the tokens and again obtain $n$ maps.
+Specifically, Smooth and Integrated CDAM are defined as follows:
+$S_{i,c}^{\text{SG}}=\frac{1}{n}\sum_{1}^{n}\sum_{j}T^{\prime}_{ij}\frac{\partial f_{c}(T^{\prime})}{\partial T^{\prime}_{ij}}$ (5)
+$S_{i,c}^{\text{IG}}=\sum_{j}(T_{ij}-\tilde{T}_{ij})\times\frac{1}{n}\sum_{1}^{n}\frac{\partial f_{c}(\tilde{T}+k/n\times(T-\tilde{T}))}{\partial T_{ij}}$ (6)
+where $T_{i}=T_{i}+\mathcal{N}\left(0,\sigma^{2}\right)$ with $\mathcal{N}\left(0,\sigma^{2}\right)$ representing Gaussian noise with standard deviation $\sigma$.
+For the baseline $\tilde{T}_{i}$ we choose the null vector.
+We implement these two variants by first propagating an image through the ViT up to the final transformer block, where the activations of the tokens are obtained.
+Next, the activations are manipulated by either applying the noise or interpolation step and the manipulated tokens are then further propagated to obtain the final ViT output.
+3.4 Quantitative Evaluation
+We perform quantitative evaluations of the correctness, class sensitivity, and compactness of importance estimators.
+In addition to our proposed methods (vanilla, smooth, and integrated CDAM), we perform evaluations on the raw attention maps, relevance propagation (RP) [21], partial Layer-wise Relevance Propagation (LRP) [22], SmoothGrad [15], Integrated Gradient [16], and Input $\times$ Gradients.
+Note that based on classic LRP [28], partial LRP [22] provides improvement for multi-head self-attention mechanisms in modern transformer architecture [22].
+Relevance propagation is a state-of-the-art importance estimator for explaining a prediction (e.g., class discriminatory) based on ViT [21].
+We use 50 interpolation steps for Integrated Gradient and Integrated CDAM and chose the standard deviation of the noise used in SmoothGrad and smooth CDAM based on the best performance on the fidelity benchmark (explained below) on a held-out validation set of 200 images.
+To evaluate the correctness (also known as faithfulness) of importance scores, we employ two versions of input perturbation-based approaches [40, 41, 42].
+Generally, an input feature (e.g., a pixel) is masked and a resulting model performance is observed.
+In all cases, input features are perturbed by being replaced with their blurred versions.
+A Gaussian blur with $\sigma=14$ turns pixels uninformative while minimizing perturbation artifacts [17].
+Fidelity.
+The first evaluation method uses accuracy curves obtained from the Most Important First (MIF) and the Least Important First (LIF) perturbations, which are abbreviated as MIF/LIF perturbation [17, 18].
+Pixels are ranked by importance scores from a specific explainability method (e.g., CDAM) and are perturbed in the MIF or LIF order.
+In the MIF perturbation, an accurate importance score should lead to a relatively rapid decrease of the model output (logit) for the target class.
+For the LIF perturbation, we expect an inverse relationship.
+Note that some importance scores are unsigned (i.e., only non-negative values) and others are signed.
+For signed importance, with negative scores faithfully indicating counter-evidence, there may be an increase in the logits if pixels with negative scores are masked.
+For unsigned and correct importance scores, one expects initially constant logits in the LIF perturbation.
+The model output $f_{\text{MIF}}(p)$ or $f_{\text{LIF}}(p)$ as a function of the fraction of perturbed pixels $p$ gives the perturbation curve [17, 18].
+We average these curves over all samples, providing an overview of fidelity for a given importance estimator.
+We further summarize the MIF and LIF perturbation evaluation by using the area $A_{\text{LIF-MIF}}$ (called the fidelity statistics) under the difference of perturbation curves $f_{\text{LIF}}-f_{\text{MIF}}$ to quantify the fidelity of the explanation methods [18].
+Generally speaking, a higher value of fidelity statistics $A_{\text{LIF-MIF}}$ indicates a better explanation [17].
+Box sensitivity.
+For the second evaluation method, inspired by Sensitivity-n [43], we measure how the model output (logit) changes as a group of pixels are perturbed.
+Pixels are selected by a randomly placed square box (of size $s\times s$); then, the sum of their importance scores and the change in the model output are measured.
+For accurate importance estimators, we expect to see a high correlation between the two measures.
+For a given image, we sample these values 100 times and calculate the Pearson correlation between the sum of importance scores and the model output change.
+This is performed for a range of differently sized boxes to obtain $f_{\text{box}}(s)$, the correlation statistics as a function of the box size $s$.
+These sensitivity curves are averaged over all samples.
+Performance of importance estimators varies with $s$, and therefore we use the area $A_{\text{box}}$ under the box sensitivity curve to quantify correctness.
+Generally, a larger box sensitivity statistics $A_{\text{box}}$ implies greater correctness.
+In contrast to Sensitivity-n which randomly selects and perturbs $n$ pixels [43], we use a random square of pixels.
+A box removes a certain region which mimics how image patches are processed by ViT.
+Box sensitivity may therefore remove larger parts or complete objects.
+In contrast, removing random pixels may leave enough neighboring pixels with sufficient semantic information such that perturbation may not have any impact of the model output.
+Indeed, we have observed that for Sensitivity-n, the correlation is very low ($<0.1$) for all methods.
+Class discrimination.
+We have designed an evaluation benchmark to measure class discrimination, which can be applied on images containing multiple objects (see $X_{\text{multi}}$ dataset below).
+We obtain explanations for two distinct classes corresponding to the present objects.
+When performing the aforementioned evaluation benchmarks (fidelity and box sensitivity), we measure the change in the model output with respect to one of the classes.
+Evaluation metrics are thus obtained using importance scores based on either the correct target class (denoted by MIF/LIF or box) or a wrong class (denoted by MIF/LIF or box).
+For important estimators that are not sensitive/discriminative to the target class, the importance scores are similar (or the same) regardless of which class is targeted.
+The change of the model output during perturbation would remain the same, irrespective of the class that was targeted to obtain the importance ranking of pixels.
+For highly class-discriminative importance estimators, importance scores would be different depending on which class is targeted.
+Furthermore, for accurate class-discriminative methods, the performances on fidelity and box sensitivity benchmarks can be expected to deteriorate substantially when perturbing according to the importance scores targeting the wrong class.
+The areas under MIF/LIF perturbation or box sensitivity curves, when using a wrong class, are denoted by $A_{\text{LIF}^{\prime}-\text{MIF}^{\prime}}$ and $A_{\text{box}^{\prime}}$, respectively.
+The areas $\Delta A_{\text{LIF-MIF}}$ and $\Delta A_{\text{box}}$ under the respective perturbation curves
+$\Delta f_{\text{LIF-MIF}}$ $=(f_{\text{LIF}}-f_{\text{MIF}})-(f_{\text{LIF}^{\prime}}-f_{\text{MIF}^{\prime}}),$ (7)
+$\Delta f_{\text{box}}$ $=f_{\text{box}}-f_{\text{box}^{\prime}}.$ (8)
+measure the degree of class-discrimination.
+In particular, if an importance estimator is not sensitive to the target class at all, $\Delta f_{\text{LIF-MIF}}\approx\Delta f_{\text{box}}\approx 0$.
+An importance estimator that performs poorly in the data randomization test of [38] would have low $\Delta A_{\text{LIF-MIF}}$ and $\Delta A_{\text{box}}$.
+Compactness.
+We evaluate sparsity and shrinkage by counting the number of input pixels with importance scores smaller than $5\%$ of the maximum value.
+A threshold is defined by $t\times S_{c}^{\text{max}}$ with $t=0.05$, where $S_{c}^{\text{max}}$ is the maximum importance score for a given image and an importance estimator.
+Generally, a compact explanation is considered desirable, where sparsity and shrinkage are quantitative approaches [44].
+3.5 Data
+We perform quantitative evaluations on two datasets.
+The first one, $X_{\text{val}}$, is a randomly selected subset of the ImageNet validation set consisting of 1000 images [19].
+The second one, $X_{\text{multi}}$, selects 1000 images from the validation set that contain multiple objects [20].
+We filtered out instances in which the class labels provided by [20] actually refer to the same object (e.g.
+notebook and laptop) by calculating and thresholding the cosine similarity of the embedded class names.
+Embeddings were obtained using OpenAI’s text-embedding-3-large model.
+The $X_{\text{multi}}$ dataset is used to evaluate the degree of class-discrimination exhibited by the explanations.
+To this end, we obtain explanations w.r.t to two annotated classes, perform the MIF/LIF perturbation and box sensitivity, and record the model output for the chosen target class.
+For example, for the top right sample in Fig.
+1, we would obtain the explanations for the classes Tench and Golden Retriever (not shown) and record the model output for the class Golden Retriever.
+For the right target (Tench), we obtain a fidelity statistics $A_{\text{LIF}-\text{MIF}}$.
+For the wrong target (Golden Retriever), $A_{\text{LIF}^{\prime}-\text{MIF}^{\prime}}$.
+In the worst case scenario, where explanations for those two distinct classes are the same, $A_{\text{LIF}^{\prime}-\text{MIF}^{\prime}}=A_{\text{LIF}-\text{MIF}}$ and the resulting $\Delta A_{\text{LIF-MIF}}$ is zero.
+Equivalent for $\Delta A_{\text{box}}$.
+For application in medical images, we use lung computed tomography (CT) scans that were collected from and validated by seven academic centers and eight medical imaging companies [23].
+The Lung Image Database Consortium image collection (LIDC) [23] contains 1018 records of CT scans.
+Presence of nodules (nodule $\leq$ 3 mm; nodule $<$ 3 mm; non-nodule $\geq$ 3 mm) is annotated by up to 4 human annotators.
+Segmentation of nodules (i.e., ROI) is the average contour given by annotators.
+Eight biomarkers that are informative in clinical practices are also collected in LIDC: subtlety, calcification, sphericity, margin, lobulation, spiculation, texture, and diameter.
+For clinical descriptions of biomarkers, refer to [45].
+After preprocessing, 854 nodule crops of size, corresponding biomarkers, and labels (benign or malignant) were saved and used for downstream procedures.
+Overall, the LIDC dataset used in this study contains 443 benign examples and 411 malignant examples.
 4 Results
-Figure 3: Class-Discriminative Attention Maps (CDAM) for user-defined concepts (Section 3.2).
-The concept embedding $l_{c}$ has been obtained by averaging the latent representations of 30 images that include the common concept.
-In each instance, from left to right: Sample image, AM, and CDAM.
-Orange and blue correspond to positive and negative values, respectively.
-For our experiments, we have been using the ViT-S/8 architecture trained with DINO as a backbone [1], which is available as a pre-trained model in their repository [^2] https://github.com/facebookresearch/dino.
-4.1 Relevance Scores from Classifier
+For our experiments using the ImageNet [19], we have been using the ViT-S/8 architecture trained with DINO as a backbone [9], which is available as a pre-trained model in their repository [^1] https://github.com/facebookresearch/dino.
+We first discuss the results of applying CDAM qualitatively and then perform a quantitative comparison with several baseline methods.
+4.1 Qualitative Evaluation
+4.1.1 CDAM for a known class
 To demonstrate our method, we have trained a linear classifier on the ImageNet [19] that achieved an accuracy of $77.0\%$.
-We use random resized cropping and horizontal flipping with PyTorch default arguments as augmentation, Adam optimizer with learning rate $3\times 10^{-4}$, batch size 128, and train for 10 epochs.
+We use random resized cropping and horizontal flipping with PyTorch default arguments as augmentation, Adam optimizer with learning rate $3\times 10^{-4}$, batch size of 128, and train for 10 epochs.
 The parameters of the ViT backbone are frozen during training, so only the classifier head is trainable.
@@
 2.
-Notice how they clearly distinguish the targeted class with positive values (evidence) from the rest of the image, which is assigned either zero (irrelevant) or negative scores (counter-evidence).
-Note that CDAM does not assign non-zero $S_{i,c}$ to tokens that have zero attention in AM.
-CDAM therefore appears to compute $S_{i,c}$ by multiplying the attention with the relevance of the corresponding token for the targeted class.
-This is a nice feature since CDAM inherits the high-quality object segmentations present in AM, which are focused on the objects of interest.
-We discuss this operating characteristic in Section 5.
-One of the most interesting features of CDAM is that it is highly sensitive to the chosen class (Fig.
-4, left).
-For example, when either the zuccini or bell pepper class is chosen, CDAM clearly highlights that vegetable.
-Most of the other vegetables are assigned either near-zero or negative values.
-Interestingly, the green stems of other vegetables are also assigned positive values.
-CDAM reveals that the ViT has learned to associate input features with chosen classes that are largely consistent with human expectations.
-Additional examples of CDAM and other interpretability methods are showned in the Appendix A (Figs.
-7, 8, 9 and 10).
-Figure 4: When targeting different classes (left) or concepts (right), the resulting CDAMs are distinct and align with the corresponding objects.
-Note that green parts of vegetables are shown to activate the zuccini class (left).
-The wrinkles in the elephant’s trunk appear to be related to the zebra concept, probably semantically mistaken by the model as zebra stripes (zoom in for a better view).
-4.2 Relevance Scores from Similarity Measure
+CDAM clearly distinguishes the region of interest (ROI) semantically related to the targeted class with positive values (yellow) from the rest of the image.
+Most backgrounds and other objects are assigned either zero (irrelevant) or negative scores (counter-evidence).
+Note that CDAM does not assign non-zero $S_{i,c}$ to tokens that have zero attention in attention maps, see also the scatter plot between CDAM and attention map Fig.
+11.
+CDAM therefore appears to compute $S_{i,c}$ by multiplying the attention with the relevance of the corresponding token for the targeted class (Section A.1).
+This is a nice feature since CDAM inherits the high-quality object segmentations present in attention maps.
+This qualitative evaluation of shrinkage and sparsity of CDAM is reflected in the quantitative measures for compactness (Section 4.2.1).
+One of the most useful features of CDAM is that it is highly sensitive to the chosen class.
+In Fig.
+4 (left), when either the zuccini or bell pepper class is chosen, CDAM clearly highlights that vegetable.
+Most of the other vegetables are assigned either near-zero or negative values (indicated in blue).
+Many other importance estimators do not distinguish the target class or only to a lesser degree.
+In Fig.
+6, targeting the burger or the hotdog does not seem to change importance scores from partial LRP, Input $\times$ Grad, and IntGrad.
+Relevance propagation demonstrates some selective focus on the ROI, but is not as discriminative as CDAM (Fig.
+6).
+This visual impression of class-sensitivity is supported by the quantitative evaluation in Section 4.2.2.
+Additional examples of CDAM and other interpretability methods are shown in the Appendix A.
+Smooth CDAM appears to reduce the regularization of importance scores, highlighting, for instance, a larger portion of the truck and cars in the background in Fig.
+5.
+This is quite different from how SmoothGrad behaves for CNNs [15], where it usually denoises the importance scores leading to a stronger focus on the ROI.
+On the other hand, Integrated CDAM can’t, for most samples, be visually distinguished from vanilla CADM even though its performance in the benchmarks is quite different (below).
+4.1.2 CDAM for a latent concept
 Concept embeddings $l_{c}$ have been obtained by averaging 30 latent representations of images depicting the concept of interest.
-These images were randomly selected from the corresponding classes in the ImageNet validation set [33] and the CDAM in Fig.
+These 30 images were randomly selected from the corresponding classes in the ImageNet validation set [19] and the CDAM in Fig.
 3 were obtained using Eq.
 4.
-Again we observe that CDAM reveals that the model clearly distinguishes the parts of the images that are relevant for the chosen concept from the rest of the image.
-For instance, for the concept hammer, CDAM highlights not only the hammer itself but also the wooden handles of the screwdrivers and the top of the ratchet, which are indeed relevant for this concept.
-Unrelated tools or parts appear in blue indicating that the model deems them to be not similar to the concept hammer.
-Just as CDAM is sensitive to the chosen class, it also produces attention maps that are specific to the chosen concept (Fig.
+We observe that CDAM reveals that the model clearly distinguishes the parts of the images that are relevant for the chosen concept from the rest of the image.
+For example, for the concept hammer (bottom right), CDAM primarily highlights the hammer.
+CDAM also provides positive importance scores for the wooden handles of the screwdrivers and the top of the ratchet, which are indeed relevant for this concept.
+Unrelated tools or parts appear in blue indicating that the model deems them to be unrelated to the concept hammer.
+Just as CDAM is sensitive to the chosen class, it also produces attention maps that are sensitive to the chosen concept (Fig.
 4, right).
-This makes the proposed method particulary valuable for understanding the representations the ViT has learned since the explanations for different concepts can give a complementary view.
-For instance, CDAM shows that parts of the elephant’s trunk are counter-evidence for the concept african elephant (Fig.
-4, right) which is at first counter-intuitive.
-However, when compared with the concept zebra, we can see that the model deems this part of the trunk to be similar to a zebra and it probably mistakes the trunk’s wrinkles for the zebra’s stripes.
-By choosing a set of images with a common concept, CDAM helps probing the latent space of the ViT, allowing us to learn about how it “sees” the world in a semi-supervised manner.
-4.3 Comparison with relevance propagation and token ablation maps
-Figure 5: Comparison of specificity to the target class.
-AM is, by definition, not influenced by it, but relevance propagation [2] does show some specificity.
-In particular due to the signed relevance scores both CDAM and TAM exhibit strong class-sensitivity and the two maps roughly agree, with CDAM appearing as stronger regularized.
-Upon visual inspection, it appears that CDAM is estimating patch token relevance that aligns well with our expectation concerning the most relevant (counter-)evidence for a given target class (Figs.
-1 and 4).
-To investigate whether these relevance scores correspond to the actual inference performed by the model, we compare them to token ablation maps (TAMs) and relevance propagation (RP) [2].
-In TAMs, the model output is measured after an individual token is removed (Section 3.3).
-RP, which is based on the Deep Taylor Decomposition, is an explanation method for ViTs that, to the best of our knowledge, produces relevance scores that are the most class-discriminative so far [2, 23].
-We find that RP is more class-discriminative than AM and that CDAM and TAM distinguish the different classes more clearly than RP ( Fig.
-5).
-Visually, the signed scores in CDAM and TAM make stronger distinctions between the different classes, whereas RP only “dims” the object that is not targeted.
-Fig.
-6 shows pair-wise comparisons among scores for CDAM, TAM, AM, and RP.
-We observe that RP tends to assign low but positive scores to tokens that obtained negative scores from CDAM or TAM.
+This makes the proposed method particularly valuable for understanding the representations the ViT has learned, since the explanations for different concepts can give a complementary view.
+For instance, CDAM distinguishes african elephant and zebra.
+However, CDAM also shows that parts of the elephant’s trunk are counter-evidence for the concept african elephant (Fig.
+4, right) which is initially counter-intuitive.
+However, when targeting the concept zebra, the model deems this part of the trunk to be similar to a zebra.
+Speculatively, the trunk may be too similar to the zebra’s muzzle or the trunk’s wrinkles may be mistaken for the zebra’s stripes.
+4.2 Quantitative evaluation
+We have performed quantitative evaluations of the correctness and compactness of our proposed methods, in comparison to multiple well-known importance estimators.
+We further investigated class sensitivity by using importance scores from multiple classes.
+4.2.1 Evaluation of correctness
 […]
-Compared to AM, TAM, and RP, CDAM appears to be less noisy (Fig.
-5).
-We can observe this implicit regularization effect by looking at the scores from 1000 randomly selected images Fig.
-6.
-In AM, TAM, and RP, the per-image average percentage of patch tokens with relevance scores smaller than $5\%$ of the maximum score are $54\%,66\%,49\%$, respectively, whereas in CDAM it is $84\%$.
-While sparsity may not always imply the most accurate estimation at a single data point (e.g., a pixel), regularization has long been pursued and motivated by interpretability in machine learning [34].
-Figure 6: Pairwise scatter plots of relevance/attention scores for CDAM, TAM, AM, and relevance propagation obtained from 1000 randomly selected images from the ImageNet validation set.
-The target class for the explanations is the one with the highest logit score.
-Notice in particular that AM and relevance propagation are highly correlated and that CDAM and TAM largely agree on the sign of relevance scores.
-We also observe that $A_{i}=0\Rightarrow S_{i,c}=0$, where $A_{i}$ and $S_{i,c}$ refer to the AM or CDAM score of token $T_{i}$.
+[…]
+4.2.2 Evaluation of class discrimination
+[…]
+[…]
+The importance scores for CDAM become anti-correlated with the model output when the wrong importance scores are used for the perturbation (Fig.
+10).
+This anti-correlation probably results from the fact that pixels that are evidence for the correct target class are counter-evidence for the wrong class.
+This result corroborates the visual impression that CDAM correctly assigns importance scores with opposite signs to the objects corresponding to the targeted class and other objects that are present (Fig.
+6).
+This is also reflected in the negative values of $f_{\text{LIF}^{\prime}}-f_{\text{MIF}^{\prime}}$ for CDAM (Fig.
+9), indicating that the ranking from least to most important has been, at least partially, reversed.
+4.2.3 Evaluation of compactness
+Sparsity and shrinkage are evaluated by our compactness evaluation, which is one of the major desired properties in explainability of deep learning [44].
+While sparsity may not always imply the most accurate estimation at a single data point, a bias-variance trade-off is well-known in machine learning [47].
+Ideally, compact explanations (e.g., sparse and shrunken importance estimators) are preferred, when other properties such as correctness and class sensitivity are constant.
+CDAM and Integrated CDAM resulted in the highest degree of compactness, together with Input $\times$ Grad.
+For those three importance estimators, on average, $88\%$ of importance scores were less than $5\%$ of the maximum score Table 2.
+IntGrad and SmoothGrad show relatively high sparsity $0.84$ and $80\%$, respectively.
+But both are lacking in correctness (Section 4.2.1) and class-discrimination (Section 4.2.2).
+Relevance Propagation shows the lowest sparsity of all considered methods $48\%$, but a high level of correctness.
+These results suggest that these correctness and compactness metrics are orthogonal.
+5 Application to Medical Images
+To demonstrate CDAM for another use case we apply it to lung nodule malignancy and biomarker prediction using the LIDC dataset [23].
+5.1 Malignancy Prediction
+After preprocessing, the LIDC data consists of 443 benign and 411 malignant lung CT scans.
+Training, validation, and test sets (in the ratios of 0.7225, 0.1275, 0.15) were stratified by and balanced according to these labels, e.g., benign and malignant.
+We fine-tuned a ViT model, with a DINO backbone pretrained on the ImageNet, for 50 epochs.
+In a parameter sweep, we varied the number of trainable layers ($10-50$) and dropout rates ($0.0-0.09$), where the learning rate was exponentially decaying ($\alpha=0.0003$ and $\beta=0.95$).
+The best accuracy on the test set of 0.85 was obtained with 50 trainable layers and dropout rate of 0.003222It might be possible to increase the model performance via further investigation.
+Here, we use a straightforward model and training scheme to demonstrate the application of CDAM on medical images..
+Attention maps and CDAM are obtained based on this model.
+The obtained attention maps suggest that the model focuses on patches with nodule fragments (Fig.
+12(b)).
+However, we note that attention maps almost always focus on pulmonary nodules without taking into account the downstream classification into benign and malignant samples.
+CDAM provides detailed structures where positive and negative values are indicated by orange and blue, respectively.
+In this binary classification where 0 is for benign and 1 for malignancy, input pixels with positive values in CDAMs are driving the classification towards malignancy.
+5.2 Biomarker prediction
+We turn our attention to clinical biomarkers in LIDC that are routinely used by medical practitioners.
+Inspired by the concept bottleneck model (CBM) [48], we fine-tuned a ViT model pre-trained with DINO on 8 biomarkers (subtlety, calcification, sphericity, margin, lobulation, spiculation, texture, and diameter).
+Essentially, a regression model was built on a pre-trained ViT model with the mean squared error (MSE) as loss function.
+For investigation of LIDC and incorporation of CBM, see [49].
+CDAMs were obtained for 8 biomarkers (Fig.
+13).
+In the context of the margin, patches including the edge of a nodule get positive CDAM scores, whereas patches corresponding to the nodule body get negative CDAM scores.
+For the diameter, patches corresponding to the nodule body get a positive score, whereas the background pixels exhibit neutral (near zero) or slightly negative scores.
+When it comes to sphericity, patches including fragments of nodule curvature get a positive CDAM score.
+Spiculation refers to the presence of needle and spike-like structures, on the margin.
+High spiculation often is an indication of malignancy.
+In a majority of cases, CDAM highlights spicules with positive values, whereas the inner bodies of nodules (i.e., not spicules) tend to get low importance scores.
 6 Discussion
-Class-Discriminative Attention Map (CDAM) is a gradient-based extension of attention maps (AM), that provides strong discrimination with respect to a chosen class or concept.
-In contrast, many post-hoc explanation methods produce similar heatmaps even when targeting different classes [23, 17].
-CDAM retains appealing characteristics of AM, namely high-quality semantic segmentation in self-supervised ViTs [1].
-Previously, a considerable effort has been made to de-noise post-hoc interpretability methods by averaging variations of multiple saliency maps [13, 15].
-Without further processing or averaging, the proposed CDAM produces comparatively sparse heat maps that are less noisy than previous methods.
-Besides explaining predictions of the classifier on the top of ViT, we also introduce how to obtain a CDAM by measuring the similarity between the latent representation of a sample image and a concept embedding.
-By probing the latent space of the ViT with different concepts (Fig.
-4) we can learn a lot about the representations it has learned.
-To some degree this can of course also be done by looking at the attention maps of the different heads, which tend to specialize in different categories of objects or parts thereof [1].
+Class-Discriminative Attention Map (CDAM) is a gradient-based extension of attention maps, that provides strong discrimination with respect to a chosen class or concept, while exhibiting excellent fidelity and compactness.
+We introduce additional variants of CDAM, inspired by SmoothGrad ([15]) and IntGrad [16]), which improve the performance on the quantitative benchmarks.
+Our proposed methods retain appealing characteristics of attention maps, namely high-quality semantic segmentation in self-supervised ViTs.
+The proposed importance estimation for CDAM scales the attention scores by how relevant they are for a target class or concept (Section A.1).
+Therefore, zero attention scores remain zero in CDAM and it produces compact explanations with high sparsity and shrinkage.
+Post-hoc explanation methods for DNNs aim to make the full decision-making process of the model more transparent by providing an approximation with certain desirable properties, such as correctness, sensitivity, and compactness.
+Particularly, many explanation methods that estimate feature importance are not very useful because they give almost identical results when targeting different classes in the model output [27, 38].
+Even randomization in the model weights is often shown to have minimal impact on explanations (i.e., saliency maps [38]).
+We find, both qualitatively and quantitatively, that CDAM is highly sensitive to the targeted class, assigning positive importance scores to objects corresponding to the target class and negative ones to other (semantically distinct) objects in the image.
+Besides explaining predictions of the classifier on the top of ViT, we introduce several variants.
+Firstly, by probing the latent space of the ViT (Fig.
+4), self-supervised CDAM provides importance scores specific to the user-defined concept.
+In the context of SSL, where class labels are absent, this makes it a valuable technique to investigate the representations learned by the model, as shared by user-selected samples.
+To some degree, this can alternatively be done by looking at the attention maps of the different heads, which tend to specialize in different categories of objects or parts thereof [9].
 This approach is, however, limited to the categories that the attention heads happened to have learned.
-CDAM allows us to probe for concepts that are relevant to any given use case.
-5.1 Proportionality of CDAM to the attention
-CDAM appears to scale the attention $A_{i}=\text{softmax}\left(\frac{Q_{\texttt{[CLS]}}K^{\top}_{i}}{\sqrt{d_{k}}}\right)$ by the relevance of the corresponding token $T_{i}$ for a concept.
-To better understand this observation, without loss of generality, consider a single attention head.
-Eq.
-4 is expanded
+Secondly, we introduce Smooth and Integrated CDAM, which essentially average a series of (vanilla) CDAMs.
+Notably, instead of adding noise to or modifying the input images as in SmoothGrad [15]) or IntGrad [16], respectively, our methods act on tokens in the final transformer block.
+Our application on medical images examplifies the need for fine-grained class-sensitive explanations offered by CDAM.
+Other explainability methods including attention maps are often highlighting whole nodules and tumors.
+But they are uninformative about the innerworkings of the ViT model predicting malignancy or biomarkers.
+While medical implication of these explanations are beyond the scope of this paper, the proof of concept application demonstrates usability of CDAM in cancer images.
+As the methods of explainable AI (XAI) becomes incorporated into practices, we foresee a greater level of human-computer interaction where class sensitivity will play a critical role for utility of explanations.
+Understanding the decision-making process of deep neural networks is intrinsically challenging due to the vast parameter space and non-linearity, which lead to complex interactions among the patch tokens in ViT.
+Post-hoc explanation methods for DNNs approximate the full operating characteristics of the model with certain desirable properties.
+Our results suggest that CDAM is capable of achieving a high degree of correctness, contrastivity, and compactness.
+Note that CDAM is not explicitly designed to satisfy these criteria.
+For example, we could impose further sparsity constraints that would increase compactness.
+However, over-optimizing an explanation method for certain properties could have undesirable effects [^3] Goodhart’s Law states that “when a measure becomes a target, it ceases to be a good measure” [50]..
+While this paper focuses on a ViT model, CDAM is applicable more broadly to all transformer models with self-attention, including large language models (LLMs) [51].
+Acknowledgments
+This work was funded by the ERA-Net CHIST-ERA grant [CHIST-ERA-19-XAI-007] long-term challenges in ICT project INFORM (ID: 93603), by the National Science Centre (NCN) of Poland [2020/02/Y/ST6/00071].
+This research was carried out with the support of the Interdisciplinary Centre for Mathematical and Computational Modelling University of Warsaw (ICM UW) under computational allocation no GDM-3540; the NVIDIA Corporation’s GPU grant; and the Google Cloud Research Innovators program.
+Appendix A Appendix
+A.1 Proportionality of CDAM to attention maps
+The proposed class-discriminative attention scores scale the attention scores.
+Empirically, in all of our evaluation tasks and applications (e.g., Fig.
+11), we observe that features with attention scores $A_{i}=\text{softmax}\left(\frac{Q_{\texttt{[CLS]}}K^{\top}_{i}}{\sqrt{d_{k}}}\right)$ of 0 have class-discriminative attention scores of 0:
+$A_{i}=0\Rightarrow S_{i,c}=0.$ (9)
+This operating characteristic is highly favorable, as besides introducing class discrimination, CDAM provides implicit regularization resulting in sparser heat maps.
+Based on a combination of empirical observations and mathematical derivations, we argue that the attention score $A_{i}=\text{softmax}\left(\frac{Q_{\texttt{[CLS]}}K^{\top}_{i}}{\sqrt{d_{k}}}\right)$ is scaled by the relevance of the corresponding token $T_{i}$ for a concept or a class.
+To better understand this observation, we first consider a concept-based CDAM with a single attention head. Eq. 4 is expanded:
 $S_{i,c}$ $=\sum_{j}T_{ij}\frac{\partial g(l,l_{c})}{\partial T_{ij}}$
@@
 […]
-Importantly, we observe in our multi-head attention ViT model (Fig. 6) that
-$A_{i}=0\Rightarrow S_{i,c}=0$ (7)
-which implies that in Section 5.1, all terms that are not proportional to $A_{i}$ are zero or cancelled out and we are left with
+Eq. 9 implies that in Section A.1, all terms that are not proportional to $A_{i}$ are zero or cancelled out and we are left with
 $S_{i,c}=A_{i}V_{i}\cdot\nabla(h\cdot l_{c}).$ (11)
-Empirically, $S_{i,c}$ is obtained by multiplying $A_{i}$ with the directional derivative of $h\cdot l_{c}$ in the direction of $V_{i}$.
+We thus find that $S_{i,c}$ is obtained by multiplying $A_{i}$ with the directional derivative of $h\cdot l_{c}$ in the direction of $V_{i}$.
 In other words, CDAM scales $A_{i}$ by the rate at which the dot product (similarity) between $h$ and $c$ changes in the direction of $V_{i}$.
-In the case that $h$ is just the identity function, we obtain an even simpler interpretation of CDAM.
-That’s because we then have $\nabla_{n}h_{k}(f)=\partial_{f_{n}}f_{k}=\delta_{nk}$ and get from Section 5.1, again assuming $\frac{\partial A_{m}}{\partial T_{ij}}=0$, that $S_{i,c}=A_{i}V_{i}\cdot l_{c}$, which means that $A_{i}$ is scaled by the similarity of $V_{i}$ and $l_{c}$.
-5.2 Faithfulness of Post-hoc Explanation Methods
-Understanding the decision-making process of deep neural networks is intrinsically challenging due to the vast parameter space and non-linearity, which lead to complex interactions among the patch tokens in ViT.
-Therefore, post-hoc explanation methods such as RP, CDAM, and TAM can only explain some facets of the model and can never be comprehensively faithful to it, as that would entail being equivalent to the model itself [17].
-Of course, the complete set of the model weights in DNNs would not be considered interpretable.
-Therefore, attempting to create explanation methods necessitates a balance among many properties, ranging from sparsity to discriminativeness [27].
-The proposed CDAM relies on gradients and is therefore a first-order approximation.
-Instead of referring directly to the input patch tokens, the CDAM scores refer to tokens that, in general, have been enriched with contributions from all other tokens through the attention layers.
-Relevance is therefore not directly attributed to input features.
-Nonetheless, by extending the inherent attention scores in ViTs, CDAM captures the high-quality semantic segmentations of input images shown in self-supervised ViTs [1], while providing class-discrimination.
-6 Conclusion
-By extending the attention maps (AM) inherent to ViT, the proposed CDAM provides relevance scores for a chosen class or concept.
-The proposed relevance estimation for CDAM essentially scales the attention scores by how relevant they are for a target class or concept.
-As a result, zero attention scores remain zero in CDAM.
-The resulting heat maps clearly distinguish the object representing the target from the background, as well as other objects that are present.
-In addition to using a classifier, we proposed to explain a user-defined concept by selecting a small set of example images.
-Using similarity measures in the latent representation, CDAM provides highly specific heat maps for a concept that underly selected examples.
-Especially in the context of SSL where class labels are absent, this makes it a valuable technique to investigate the model representations.
-In terms of class-discrimination, our method compares favorably against the state of the art, namely relevance propagation (RP) for transformers [2].
-Especially, the positive and negative signs of relevance scores in CDAM make clear distinctions qualitatively and quantitatively.
-We also devise the token ablation maps (TAM) which visualize the changes in the ViT model output when corresponding image patch tokens are removed.
-The advantage of TAM is its explicable computation, which serves as a sanity check.
-CDAM captures the overall patterns of TAM while providing sparser and more semantically consistent heat maps.
-However, while RP and TAM are used in comparison, neither of them is the “ground truth” of relevance scores.
-Post-hoc explanation methods for DNNs approximate the full operating characteristics of the model with certain desirable properties.
-Our results suggest that CDAM is capable of achieving a high degree of correctness, contrastivity, and compactness, as described in [35].
-Note that CDAM is not explicitly designed to satisfy these criteria.
-For example, we could impose further sparsity constraints that would increase compactness.
-However, over-optimizing an explanation method for certain properties could have undesirable effects [^3] Goodhart’s Law states that “when a measure becomes a target, it ceases to be a good measure” [36].
-While this paper uses a ViT model trained with SSL [1], CDAM is applicable more broadly to all transformer models with self-attention, e.g., large language models [37].
-Acknowledgments
-This work was funded by the ERA-Net CHIST-ERA grant [CHIST-ERA-19-XAI-007] long term challenges in ICT project INFORM (ID: 93603), by the National Science Centre (NCN) of Poland [2020/02/Y/ST6/00071].
-This research was carried out with the support of the Interdisciplinary Centre for Mathematical and Computational Modelling University of Warsaw (ICM UW) under computational allocation no GDM-3540; the NVIDIA Corporation’s GPU grant; and the Google Cloud Research Innovators program.
-Appendix A Appendix
-A.1 Multiple attention heads
-We note that introducing $h$ attention heads does not change the derivations in Section 5.1 qualitatively.
-Instead of the sum $Z=\sum_{m}A_{m}V_{m}$ there will be a term $\text{Concat}(Z^{1},Z^{2},..,Z^{h})W^{O},W^{O}\in\mathbb{R}^{hd_{v}\times d_{\text{model}}}$ so that in the case of multi-head attention the equivalent of Section 5.1 reads
+In the special case that $h$ is the identity function, we obtain an even simpler interpretation of CDAM due to $\nabla_{n}h_{k}(f)=\partial_{f_{n}}f_{k}=\delta_{nk}$.
+Assuming $\frac{\partial A_{m}}{\partial T_{ij}}=0$, Section A.1 simplifies to $S_{i,c}=A_{i}V_{i}\cdot l_{c}$, which means that $A_{i}$ is scaled by the similarity of $V_{i}$ and $l_{c}$.
+We note that introducing $h$ attention heads does not change the above derivations qualitatively.
+Instead of the sum $Z=\sum_{m}A_{m}V_{m}$, there will be a term $\text{Concat}(Z^{1},Z^{2},..,Z^{h})W^{O},W^{O}\in\mathbb{R}^{hd_{v}\times d_{\text{model}}}$ so that in the case of multi-head attention the equivalent of Section A.1 reads
 $S_{i,c}=\text{Concat}\left(Y^{1},Y^{2},..,Y^{h}\right)W^{O}\nabla(h\cdot l_{c}),\;Y^{n}=T_{i}\sum_{m}\frac{\partial A^{n}_{m}}{\partial T_{i}}V^{n}_{m}+A^{n}_{i}V^{n}_{i}$ (12)
@@
 Each of the summands is proportional to the corresponding $A^{n}$, and the interpretation of CDAM from the single-head case therefore extends to the multi-head one.
-Although our observation $A_{i}=0\Rightarrow S_{i,c}=0$ implies that all terms that are not proportional to $A_{i}$ in Section 5.1 should vanish, we don’t have a theoretical argument why that would be the case.
+Although our observation $A_{i}=0\Rightarrow S_{i,c}=0$ implies that all terms that are not proportional to $A_{i}$ in Section A.1 should vanish, we don’t have a theoretical argument why that would be the case.
 It seems unlikely that the various terms in the first sum cancel out, which leaves $\frac{\partial A_{m}}{\partial T_{ij}}=0$ as an explanation.
 However, in general, $A_{i}$ depends on $T_{i}$ through $K_{i}$ and one would therefore expect the derivative to be non-zero.
-We can also show that CDAM scales the attention by the relevance of the correpsonding token for a chosen class (instead of concept), in the case that a classifier with weights $W^{C}$ has been trained on top of the ViT.
+We can also show that CDAM scales the attention by the relevance of the corresponding token for a chosen class (instead of concept), in the case that a classifier with weights $W^{C}$ has been trained on top of the ViT.
 Analogously to the above single-head discussion we have
@@
 All the intuitions we have gained from the concept-based CDAM therefore extend to the class-based one.
-A.2 Examples for explanation methods
-Figure 7: Examples for the considered explanation methods. The indicated class is the one predicted by the model and is the target for the explanation methods.
-Figure 8: Examples for the considered explanation methods. The indicated class is the one predicted by the model and is the target for the explanation methods.
-Figure 9: Examples for the considered explanation methods. The indicated class is the one predicted by the model and is the target for the explanation methods.
-Figure 10: Examples for the considered explanation methods. The indicated class is the one predicted by the model and is the target for the explanation methods.
+A.2 Results for a random subset of ImageNet
+Figure S1: Box sensitivity curves obtained from the random ImageNet subset
+Figure S2: MIF/LIF perturbation curves obtained from the random ImageNet subset.
+[…]
+CDAM 801 1423 623 11.3
+Integrated CDAM 696 1459 753 12.9
+Smooth CDAM 591 1482 890 11.9
+Rel. Prop. 475 1432 957 12.7
+Partial LRP 654 1329 674 8.4
+Att. map 580 1352 771 9.1
+SmoothGrad 796 786 -10 0.2
+IntGrad 411 1055 644 8.3
+Input x Grad 470 935 465 3.1
+Table S1: Area under the curve for MIF, LIF and LIF-MIF curves (Fig. S2) and box sensitivity curves (Fig. S1) obtained from the random ImageNet subset.
+A.3 Examples of CDAM for a known class on the ImageNet
+Figure S3: Examples for CDAM and other explanation methods on the ImageNet. The indicated class is the one predicted by the model and is the target for the explanation methods.
+Figure S4: Examples for CDAM and other explanation methods on the ImageNet. The indicated class is the one predicted by the model and is the target for the explanation methods.
+Figure S5: Examples for CDAM and other explanation methods on the ImageNet. The indicated class is the one predicted by the model and is the target for the explanation methods.
+A.4 Examples for CDAM on the Lung Image Database Consortium (LIDC)
+Figure S6: Examples for the Attention Map and CDAM based on a ViT malignancy model trained on the lung nodule data from LIDC.
+Figure S7: Examples for the Attention Map and CDAM based on a ViT biomarker model trained on the lung nodule data from LIDC.
+Figure S8: Examples for the Attention Map and CDAM based on a ViT biomarker model trained on the lung nodule data from LIDC.

### v3

@@
 Importance estimators are explainability methods that quantify feature importance for deep neural networks (DNN).
-In vision transformers (ViT), the self-attention mechanism naturally leads to attention maps, which are sometimes used as importance scores for which input features ViT models are focusing on.
+In vision transformers (ViT), the self-attention mechanism naturally leads to attention maps, which are sometimes interpreted as importance scores that indicate which input features ViT models are focusing on.
 However, attention maps do not account for signals from downstream tasks.
@@
 Additionally, we introduce Smooth CDAM and Integrated CDAM, which average a series of CDAMs with slightly altered tokens.
-Our quantitative benchmarks include correctness, compactness, and class sensitivity, in comparison to six other importance estimators.
+Our quantitative benchmarks include correctness, compactness, and class sensitivity, in comparison to 7 other importance estimators.
 Vanilla, Smooth, and Integrated CDAM excel across all three benchmarks.
@@
 Overall, CDAM is shown to be highly class-discriminative and semantically relevant, while providing compact explanations.
-PyTorch implementation: https://github.com/lenbrocki/CDAM.
-Web live demo: https://cdam.informatism.com/.
+Figure 1: Extending attention maps (AM), the proposed class-discriminative attention maps (CDAM) quantify and visualize input features that are relevant for the target class in Vision Transformers (ViT).
+We visualize importance scores obtained with CDAM for a linear classifier with ViT-S/8, trained with DINO, as a backbone (Caron et al., 2021).
+For details, refer to Section 3.1.
+Orange and blue colors correspond to positive and negative values, respectively.
 1 Introduction
-Vision transformers (ViT) are a family of computer vision models based on the transformer architecture [1], that demonstrate comparable or even better performance than convolutional neural networks (CNN) on many object detection and classification tasks [2, 3].
-A powerful feature of the transformers is that the attention mechanism provides a built-in explanation of the model’s inner workings, if attention is understood as a weighing of the relevance of individual tokens [4, 5, 6, 7].
+Vision transformers (ViT) are a family of computer vision models based on the transformer architecture (Vaswani et al., 2017), that demonstrate comparable or even better performance than convolutional neural networks (CNN) on many object detection and classification tasks (Dosovitskiy et al., 2020; Wu et al., 2020).
+A powerful feature of the transformers is that the attention mechanism provides a built-in explanation of the model’s inner workings, if attention is understood as a weighing of the relevance of individual tokens (Mullenbach et al., 2018; Bahdanau et al., 2014; Serrano & Smith, 2019; Thorne et al., 2019).
 While attention maps provide intuitive and inherent interpretability, they fail to be discriminative with respect to a target class.
-Recognizing the potential of attention maps for explaining a ViT’s inner workings, we have developed class-discriminative attention maps (CDAM), a gradient-based extension that makes them sensitive to class-specific signals based on a classifier or a concept derived from selected samples.
+Recognizing the potential of attention maps for explaining a ViT’s inner workings, we have developed class-discriminative attention maps (CDAM), a gradient-based extension that makes them sensitive to class-specific signals based on a classifier or a latent concept.
 Attention maps are the visualization of the self-attention of the class token query [CLS] in the final attention layer.
-Attention heads can be visualized separately, and in certain instances, may be related to semantically meaningful concepts (e.g., in language [8]).
-Recently, ViT trained with self-supervised learning (SSL) methods such as DINO [9] have shown excellent performance leveraging a greater amount of unlabeled data.
+Attention heads can be visualized separately, and in certain instances, may be related to semantically meaningful concepts (e.g., in language (Clark et al., 2019)).
+Recently, ViT trained with self-supervised learning (SSL) methods (Caron et al., 2021; Touvron et al., 2021; Oquab et al., 2024) have shown excellent performance leveraging a greater amount of unlabeled data.
 Attention maps from such self-supervised ViT models provide high-quality semantic segmentations of input images.
-These attention maps clearly distinguish regions of interest (ROI) from the background with very little noise, with different attention heads often focusing on semantically distinct regions [9].
-There have been many improvements to ViT including the introduction of convolution [10, 11] and spatial sparsity [12].
+These attention maps clearly distinguish regions of interest (ROI) from the background with very little noise, with different attention heads often focusing on semantically distinct regions (Caron et al., 2021).
+There have been many improvements to ViT including the introduction of convolution (Wu et al., 2021; Yuan et al., 2021) and spatial sparsity (Li et al., 2021).
 Attention maps, as well as our proposed CDAMs, are applicable to all ViT architectures with self-attention.
@@
 Therefore, CDAM is a useful tool for explaining the decision-making process of classifiers with a ViT backbone.
+Note that the goal of CDAM is to explain the classifier head and the final transformer block.
 In addition to the classifier-based explanations, we present a concept-based approach that replaces the class score with the similarity measure in the latent space of the ViT, where a concept is defined through selected images.
-Averaging the latent representations of the example images yields a concept embedding, which is analogous to a concept vector in variational autoencoders (VAE) and related generative models [13, 14].
+Averaging the latent representations of the example images yields a concept embedding, which is analogous to a concept vector in variational autoencoders (VAE) and related generative models (Brocki & Chung, 2019).
 This allows obtaining a heat map for the target concept the model has never been explicitly trained on, which is of particular interest for self-supervised models.
 We further demonstrate when computing gradients of a class (Section 3.1) or concept similarity (Section 3.2) score, with respect to the token activations in the final transformer block, smoothing or averaging could be applied to improve performance.
-Specifically, we introduce two additional variants of CDAM; (1) Smooth CDAM averages over multiple gradients that are computed with additional noise to tokens (inspired by SmoothGrad [15]) and (2) Integrated CDAM takes the intergral of the gradients along the path from the baseline to the token in the final transformer block (inspired by IntGrad [16]).
+Specifically, we introduce two additional variants of CDAM; (1) Smooth CDAM averages over multiple gradients that are computed with additional noise to tokens (inspired by SmoothGrad (Smilkov et al., 2017)) and (2) Integrated CDAM takes the integral of the gradients along the path from the baseline to the token in the final transformer block (inspired by IntGrad (Sundararajan et al., 2017)).
 We conduct several quantitative evaluations focusing on correctness, class sensitivity, and compactness.
-Correctness is measured by perturbing an increasing amount of input features and measuring the model output [17, 18].
-By using the ImageNet samples [19] with multiple objects [20] and applying importance estimators for different classes, we quantify the level of class-discrimination.
+Correctness is measured by perturbing an increasing amount of input features and measuring the model output (Brocki & Chung, 2023c; b).
+By using the ImageNet samples (Deng et al., 2009) with multiple objects (Beyer et al., 2020) and applying importance estimators for different classes, we quantify the level of class-discrimination.
 Sparsity and shrinkage of explanations are measured by the number of near-zero importance scores.
-Besides vanilla, Smooth, and Integrated CDAM, we also employed attention maps, relevance propagation (RP) [21], partial Layer-wise Relevance Propagation (LRP) [22], Input $\times$ Gradients, SmoothGrad [15], and Integrated Gradient [16] in comparison.
-Three variants of CDAM outperform other explainability methods on three benchmarks.
-Lastly, we have applied CDAM on a ViT fine-tuned on the Lung Image Database Consortium image collection (LIDC) [23].
+Besides the proposed vanilla, Smooth, and Integrated CDAM, we also employed 7 different explainability methods in comparison.
+Three variants of CDAM outperform other methods on quantitative benchmarks.
+Lastly, we have applied CDAM on a ViT fine-tuned on the Lung Image Database Consortium image collection (LIDC) (Armato III et al., 2011).
 Two ViT-based models are built for predicting malignancy and biomarkers.
-CDAM, as well as attention maps, are obtained.
 CDAM is shown to be highly class-discriminative and semantically relevant, while providing implicit regularization of importance scores.
@@
 Gradient-based methods operate by backpropagating gradients from the prediction score to the input features to produce feature attribution maps, which is also called saliency maps.
-Many variants have been proposed, see [24, 16, 25, 15, 26] to just name a few.
-In particular, CDAM is similar to and inspired by the input×$\times$×gradient method [26], because it involves calculating the gradients with respect to tokens multiplied by their activation (Eqs.
+Many variants have been proposed, see (Simonyan et al., 2013; Sundararajan et al., 2017; Selvaraju et al., 2017; Smilkov et al., 2017; Shrikumar et al., 2016) to just name a few.
+In particular, CDAM is similar to and inspired by the input×$\times$×gradient method (Shrikumar et al., 2016), because it involves calculating the gradients with respect to tokens multiplied by their activation (Eqs.
 2 and 4).
@@
 CDAM stops backpropagation at the tokens that enter the final attention layer.
-In that regard, our method is related to GradCam [25] which backpropagates the gradients to the final convolutional feature map in a CNN.
-GradCam and CDAM therefore share the property that they operate on high-level features which is presumably the reason why they are both sensitive to the targeted class, a feature that many explanation methods do not have [27].
-Several methods have attempted to apply gradient-based methods or Layer-Wise Relevance Propagation (LRP) [28] to Transformer architectures [29, 30, 22, 31, 32], mostly in the context of NLP tasks.
-Attention has often been used to gain insight into a model’s inner workings [4, 5, 6, 7] since it seems intuitive that this weighing of individual tokens correlates with how important they are for the performed task.
-There is also an ongoing discussion about whether attention provides meaningful explanations [6, 33, 34].
-Furthermore, methods such as “attention rollout” [31] have been proposed that aim to quantify how the attention flows through the model to the input tokens.
-[21] defines the relevancy matrices, as the Hadamard product of the attention map and the gradient of $f_{c}$ w.r.t.
+In that regard, our method is related to GradCam (Selvaraju et al., 2017) which backpropagates the gradients to the final convolutional feature map in a CNN.
+GradCam and CDAM therefore share the property that they operate on high-level features which is presumably the reason why they are both sensitive to the targeted class, a feature that many explanation methods do not have (Rudin, 2019).
+Several methods have attempted to apply gradient-based methods or Layer-Wise Relevance Propagation (LRP) (Bach et al., 2015) to Transformer architectures (Atanasova et al., 2020; Ali et al., 2022; Voita et al., 2019; Abnar & Zuidema, 2020; Chefer et al., 2021b), mostly in the context of NLP tasks.
+GradCam for ViT backpropages the gradients to the outputs of the final attention layer (Gildenblat, 2021).
+Attention has often been used to gain insight into a model’s inner workings (Mullenbach et al., 2018; Bahdanau et al., 2014; Serrano & Smith, 2019; Thorne et al., 2019) since it seems intuitive that this weighing of individual tokens correlates with how important they are for the performed task.
+There is also an ongoing discussion about whether attention provides meaningful explanations (Serrano & Smith, 2019; Wiegreffe & Pinter, 2019; Jain & Wallace, 2019).
+Furthermore, methods such as “attention rollout” (Abnar & Zuidema, 2020) have been proposed that aim to quantify how the attention flows through the model to the input tokens.
+(Chefer et al., 2021a) defines the relevancy matrices, as the Hadamard product of the attention map and the gradient of $f_{c}$ w.r.t.
 to the attention layer, accumulated at each layer.
-This relevance propagation (RP) method [21], to the best of our knowledge, shows the highest degree of class-discrimination in ViT models.
-This is used in our study as one of comparative importance estimators.
-Many papers have been published on the problem of evaluating post-hoc explanation methods, leading even to the proposal of meta-evaluations [35].
-In a recent overview of quantitative evaluation methods [36] 12 desirable properties have been identified.
+This relevance propagation (RP) method (Chefer et al., 2021a), to the best of our knowledge, shows the highest degree of class-discrimination in ViT models.
+This is used in our study as one of the baseline importance estimators.
+Many papers have been published on the problem of evaluating post-hoc explanation methods, leading even to the proposal of meta-evaluations (Hedström et al., 2023).
+In a recent overview of quantitative evaluation methods (Nauta et al., 2023b) 12 desirable properties have been identified.
 We primarily focus on three of them, namely correctness, contrastivity, and compactness.
-In particular, carefully designed occulusion/perturbation studies can help us calculate a fidelity statistics [17], which is also called the symmetric relevance gain (SRG) measure [37].
-[38] first demonstrated that many importance estimators, which purport to be sensitive to the internals of a classifier, produce similar importance scores regardless of the relation of inputs and classes in the training data, which they called “the data randomization test”.
-Generally, if two importance scores with respect to two different classes produce very similar heat maps, the importance estimator is not class sensitive.
+In particular, carefully designed occlusion/perturbation studies can help us calculate the fidelity statistics (Brocki & Chung, 2023c), which is also called the symmetric relevance gain (SRG) measure (Bluecher et al., 2024).
+Adebayo et al.
+(2018) first demonstrated that many importance estimators, which purport to be sensitive to the internals of a classifier, produce similar importance scores regardless of the relation of inputs and classes in the training data, which they called “the data randomization test”.
+Generally, if two importance scores with respect to two different classes produce very similar heat maps, the importance estimator is not class-sensitive.
 We quantify this concept in our class sensitivity evaluation benchmark to measure whether an explanation is sensitive to the classes.
 3 Methods
-Adaptation of the transformer architecture to computer vision hinges on creating tokens from small patches of the input image [2].
-In addition to those image patch tokens, the [CLS] token is added, which is utilized for downstream tasks such as classification.
-Attention maps are the two-dimensional visualization of the self-attention in the final attention layer with [CLS] as query token, where the attention scores are averaged over all heads.
+[…]
+Figure 2: Graphical scheme of class-discriminative attention maps (CDAM).
+(a) CDAMs are obtained by first propagating an image through the transformer blocks and using its latent representation $l=\texttt{[CLS]}^{\prime}$ to infer a class score $f_{c}$ or similarity score $g(l,l_{c})$ (see Eq.
+2 or Eq.
+4).
+(b) Detailed view of the final transformer block, adapted from (Dosovitskiy et al., 2020).
 […]
@@
 by how relevant the corresponding tokens $T_{i}$ are for the class or similarity score.
-$Q$ and $K$ are the query and key matrices, consisting of a set of $d_{k}$ dimensional vectors [1].
-We discuss the proportionality of CDAM to the attention scores in Section A.1.
-Self-supervised learning (SSL) methods such as DINO can train ViT models whose attention maps are high-quality semantic segmentations of the input, without using labels [9].
-Therefore, we utilize a ViT model [Dosovitskiy2020ViT] pre-trained with DINO as a backbone for our experiments and applications.
-We denote details about the ViT architecture, training schemes, and model performance in the Section 4.
+Here $Q_{\texttt{[CLS]}}=T_{\texttt{[CLS]}}W^{Q}$ with $W^{Q}\in\mathbb{R}^{d_{\text{model }}\times d_{k}}$ and $K_{i}=T_{i}W^{K}$ with $W^{K}\in\mathbb{R}^{d_{\text{model }}\times d_{k}}$.
+In this notation $Q$ and $K$ are the query and key matrices which pack together all the query and key vectors (Vaswani et al., 2017) and the indices, e.g.
+$K_{i}$, select single query or key vectors.
+Self-supervised learning (SSL) methods such as DINO can train ViT models whose attention maps are high-quality semantic segmentations of the input, without using labels (Caron et al., 2021).
+Therefore, we utilize a ViT model (Dosovitskiy et al., 2020) pre-trained with DINO as a backbone for our main experiments and applications.
+Furthermore, we demonstrate general applicability of CDAM by using other ViT models, namely pre-trained with Supervised Weakly through Hashtags (SWAG) (Singh et al., 2022), Data-efficient image Transformers (DeIT) (Touvron et al., 2021) and DINOv2 (Oquab et al., 2024).
+We denote details about the ViT architecture, training schemes, and model performance in the Section 4 and Section A.5.
 Note that the heat maps shown in this article are scaled independently.
@@
 2).
-We found in our experiments that calculating the gradients with respect to the layer-normalized tokens yields far better results than using the unnormalized ones.
+We found in our experiments that calculating the gradients with respect to the layer-normalized tokens yields far better results than using the not normalized ones.
 This definition is a first-order estimation of the contribution of $T_{i}$ to $f_{c}$ in the sense that it would be exact if $f_{c}$ depended on $T_{ij}$ only linearly.
-It is inspired by gradient-based explanation methods with the difference that we don’t backpropagate the gradients all the way to the input layer but only to the input tokens to the final attention layer.
+It is inspired by gradient-based explanation methods with the difference that we don’t backpropagate the gradients all the way to the input layer but only to the input tokens of the final attention layer.
 This allows us to gain insight into the decision-making process of the combination of ViT and classifier based on high-level features.
@@
 Another way of probing high-level features learned by the ViT is through the use of a similarity measure $g$ in the latent space of the ViT.
-We compare the latent representation $l(X)=T_{\texttt{[CLS]}^{\prime}}(X)$ of a sample image $X$ and a concept vector $l_{c}$ [39, 13, 14], i.e.
+We compare the latent representation $l(X)=T_{\texttt{[CLS]}^{\prime}}(X)$ of a sample image $X$ and a concept vector $l_{c}$ (Brocki & Chung, 2019), i.e.
 the averaged latent representation of $n$ images that are chosen to represent the concept of interest:
 $l_{c}=\frac{1}{n}\sum_{X\in c}l(X).$ (3)
-This approach is directly inspired by the concept saliency map [13]. The definition of the importance score in this setting is analogous to Eq. 2
+This approach is directly inspired by the concept saliency map (Brocki & Chung, 2019). The definition of the importance score in this setting is analogous to Eq. 2
 $S_{i,c}=\sum_{j}T_{ij}\frac{\partial g(l,l_{c})}{\partial T_{ij}}.$ (4)
@@
 3.3 Smooth and Integrated CDAM
-SmoothGrad (SG) [15] and Integrated Gradients (IG) [16] extended saliency maps [24], which were introduced in the context of CNNs as visualizing gradients of the class with respect to the input features.
-Essentially, these two modifications sharpen and denoise vanilla saliency maps, with theoretical and practical justifications.
-Briefly, [15] adds a small Gaussian noise to a sample image $n$ times, resulting in $n$ vanilla saliency maps.
+SmoothGrad (SG) (Smilkov et al., 2017) and Integrated Gradients (IG) (Sundararajan et al., 2017) extended saliency maps (Simonyan et al., 2013), which were introduced in the context of CNNs as visualizing gradients of the class with respect to the input features.
+Briefly, (Smilkov et al., 2017) adds a small Gaussian noise to a sample image $n$ times, resulting in $n$ vanilla saliency maps.
 SG is an average of those $n$ saliency maps.
-In IG, instead of adding random noise, [16] uses multiple saliency maps obtained from interpolating between a given sample and a baseline (e.g., a black image).
+In IG, instead of adding random noise, (Sundararajan et al., 2017) uses multiple saliency maps obtained from interpolating between a given sample and a baseline (e.g., a black image).
 We have adapted the analogous modifications to CDAM, resulting in two additional variants which we call Smooth and Integrated CDAM.
-Instead of modifying the input images, we vary the tokens that enter the final transformer block.
+Importantly, instead of modifying the input images, we vary the tokens that enter the final transformer block.
 In Integrated CDAM, we interpolate between a baseline and the tokens themselves, resulting in $n$ maps.
-In Smooth CDAM, we repeatedly add a small noise to the tokens and again obtain $n$ maps.
+Notice that, since we apply the same method, the axioms of IG still hold with the only difference that they apply to the tokens that enter the final transformer block instead of the input pixels.
+In Smooth CDAM, we repeatedly add a small Gaussian noise to the tokens and again obtain $n$ maps.
 Specifically, Smooth and Integrated CDAM are defined as follows:
@@
 $S_{i,c}^{\text{IG}}=\sum_{j}(T_{ij}-\tilde{T}_{ij})\times\frac{1}{n}\sum_{1}^{n}\frac{\partial f_{c}(\tilde{T}+k/n\times(T-\tilde{T}))}{\partial T_{ij}}$ (6)
-where $T_{i}=T_{i}+\mathcal{N}\left(0,\sigma^{2}\right)$ with $\mathcal{N}\left(0,\sigma^{2}\right)$ representing Gaussian noise with standard deviation $\sigma$.
-For the baseline $\tilde{T}_{i}$ we choose the null vector.
+where $T^{\prime}_{i}=T_{i}+\mathcal{N}\left(0,\sigma^{2}\right)$ with $\mathcal{N}\left(0,\sigma^{2}\right)$ representing Gaussian noise with standard deviation $\sigma$.
+For our experiments, we used $n=50$.
+For the baseline $\tilde{T}_{i}$ we choose the null vector of 0.
 We implement these two variants by first propagating an image through the ViT up to the final transformer block, where the activations of the tokens are obtained.
@@
 We perform quantitative evaluations of the correctness, class sensitivity, and compactness of importance estimators.
-In addition to our proposed methods (vanilla, smooth, and integrated CDAM), we perform evaluations on the raw attention maps, relevance propagation (RP) [21], partial Layer-wise Relevance Propagation (LRP) [22], SmoothGrad [15], Integrated Gradient [16], and Input $\times$ Gradients.
-Note that based on classic LRP [28], partial LRP [22] provides improvement for multi-head self-attention mechanisms in modern transformer architecture [22].
-Relevance propagation is a state-of-the-art importance estimator for explaining a prediction (e.g., class discriminatory) based on ViT [21].
-We use 50 interpolation steps for Integrated Gradient and Integrated CDAM and chose the standard deviation of the noise used in SmoothGrad and smooth CDAM based on the best performance on the fidelity benchmark (explained below) on a held-out validation set of 200 images.
-To evaluate the correctness (also known as faithfulness) of importance scores, we employ two versions of input perturbation-based approaches [40, 41, 42].
+In addition to our proposed methods (vanilla, Smooth, and Integrated CDAM), we perform evaluations on attention maps, Input $\times$ Gradients, relevance propagation (RP) (Chefer et al., 2021a), partial Layer-wise Relevance Propagation (LRP) (Voita et al., 2019), SmoothGrad (Smilkov et al., 2017), IntGrad (Sundararajan et al., 2017), and GradCam for ViT (Gildenblat, 2021).
+Note that based on classic LRP (Bach et al., 2015), partial LRP (Voita et al., 2019) provides improvement for multi-head self-attention mechanisms in modern transformer architecture (Voita et al., 2019).
+Relevance propagation is a state-of-the-art importance estimator for explaining a prediction (e.g., class discriminatory) based on ViT (Chefer et al., 2021a).
+For IntGrad, SmoothGrad, Integrated CDAM, and Smooth CDAM, $n=50$ is used.
+The noise ($\sigma$) used in SmoothGrad and Smooth CDAM is selected based on the best performance on a validation set of 200 images.
+To evaluate the correctness (also known as faithfulness) of importance scores, we employ two versions of input perturbation-based approaches (Samek et al., 2016; Petsiuk et al., 2018; Kindermans et al., 2017).
 Generally, an input feature (e.g., a pixel) is masked and a resulting model performance is observed.
 In all cases, input features are perturbed by being replaced with their blurred versions.
-A Gaussian blur with $\sigma=14$ turns pixels uninformative while minimizing perturbation artifacts [17].
+A Gaussian blur with $\sigma=14$ turns pixels uninformative while minimizing perturbation artifacts (Brocki & Chung, 2023c).
 Fidelity.
-The first evaluation method uses accuracy curves obtained from the Most Important First (MIF) and the Least Important First (LIF) perturbations, which are abbreviated as MIF/LIF perturbation [17, 18].
+The first evaluation method uses accuracy curves obtained from the Most Important First (MIF) and the Least Important First (LIF) perturbations, which are abbreviated as MIF/LIF perturbation (Brocki & Chung, 2023c; b).
 Pixels are ranked by importance scores from a specific explainability method (e.g., CDAM) and are perturbed in the MIF or LIF order.
@@
 For unsigned and correct importance scores, one expects initially constant logits in the LIF perturbation.
-The model output $f_{\text{MIF}}(p)$ or $f_{\text{LIF}}(p)$ as a function of the fraction of perturbed pixels $p$ gives the perturbation curve [17, 18].
+The model output $f_{\text{MIF}}(p)$ or $f_{\text{LIF}}(p)$ as a function of the fraction of perturbed pixels $p$ gives the perturbation curve (Brocki & Chung, 2023c; b).
 We average these curves over all samples, providing an overview of fidelity for a given importance estimator.
-We further summarize the MIF and LIF perturbation evaluation by using the area $A_{\text{LIF-MIF}}$ (called the fidelity statistics) under the difference of perturbation curves $f_{\text{LIF}}-f_{\text{MIF}}$ to quantify the fidelity of the explanation methods [18].
-Generally speaking, a higher value of fidelity statistics $A_{\text{LIF-MIF}}$ indicates a better explanation [17].
+We further summarize the MIF and LIF perturbation evaluation by using the area $A_{\text{LIF-MIF}}$ (called the fidelity statistics) under the difference of perturbation curves $f_{\text{LIF}}-f_{\text{MIF}}$ to quantify the fidelity of the explanation methods (Brocki & Chung, 2023b).
+Generally speaking, a higher value of fidelity statistics $A_{\text{LIF-MIF}}$ indicates a better explanation (Brocki & Chung, 2023c).
+Similar ideas are developed into the symmetric relevance gain (SRG) measure (Bluecher et al., 2024).
 Box sensitivity.
-For the second evaluation method, inspired by Sensitivity-n [43], we measure how the model output (logit) changes as a group of pixels are perturbed.
+For the second evaluation method, inspired by Sensitivity-n (Ancona et al., 2017), we measure how the model output (logit) changes as a group of pixels is perturbed.
 Pixels are selected by a randomly placed square box (of size $s\times s$); then, the sum of their importance scores and the change in the model output are measured.
@@
 Generally, a larger box sensitivity statistics $A_{\text{box}}$ implies greater correctness.
-In contrast to Sensitivity-n which randomly selects and perturbs $n$ pixels [43], we use a random square of pixels.
+In contrast to Sensitivity-n which randomly selects and perturbs $n$ pixels (Ancona et al., 2017), we use a random square of pixels.
 A box removes a certain region which mimics how image patches are processed by ViT.
@@
 In particular, if an importance estimator is not sensitive to the target class at all, $\Delta f_{\text{LIF-MIF}}\approx\Delta f_{\text{box}}\approx 0$.
-An importance estimator that performs poorly in the data randomization test of [38] would have low $\Delta A_{\text{LIF-MIF}}$ and $\Delta A_{\text{box}}$.
+An importance estimator that performs poorly in the data randomization test of (Adebayo et al., 2018) would have low $\Delta A_{\text{LIF-MIF}}$ and $\Delta A_{\text{box}}$.
 Compactness.
@@
 A threshold is defined by $t\times S_{c}^{\text{max}}$ with $t=0.05$, where $S_{c}^{\text{max}}$ is the maximum importance score for a given image and an importance estimator.
-Generally, a compact explanation is considered desirable, where sparsity and shrinkage are quantitative approaches [44].
+Generally, a compact explanation is considered desirable, where sparsity and shrinkage are quantitative approaches (Nauta et al., 2023a).
 3.5 Data
-We perform quantitative evaluations on two datasets.
-The first one, $X_{\text{val}}$, is a randomly selected subset of the ImageNet validation set consisting of 1000 images [19].
-The second one, $X_{\text{multi}}$, selects 1000 images from the validation set that contain multiple objects [20].
-We filtered out instances in which the class labels provided by [20] actually refer to the same object (e.g.
+We have used two subsets of the ImageNet (denoted as $X_{\text{val}}$ and $X_{\text{multi}}$) (Deng et al., 2009), as well as lung computed tomography (CT) scans from the Lung Image Database Consortium image collection (LIDC) (Armato III et al., 2011).
+The first one, $X_{\text{val}}$, is a randomly selected subset of the ImageNet validation set consisting of 1000 images (Deng et al., 2009).
+The second one, $X_{\text{multi}}$, selects 1000 images from the validation set that contain multiple objects (Beyer et al., 2020).
+We filtered out instances in which the class labels provided by (Beyer et al., 2020) actually refer to the same object (e.g.
 notebook and laptop) by calculating and thresholding the cosine similarity of the embedded class names.
@@
 Equivalent for $\Delta A_{\text{box}}$.
-For application in medical images, we use lung computed tomography (CT) scans that were collected from and validated by seven academic centers and eight medical imaging companies [23].
-The Lung Image Database Consortium image collection (LIDC) [23] contains 1018 records of CT scans.
+The Lung Image Database Consortium image collection (LIDC) contains 1018 records of lung CT scans, that were collected from and validated by seven academic centers and eight medical imaging companies (Armato III et al., 2011).
 Presence of nodules (nodule $\leq$ 3 mm; nodule $<$ 3 mm; non-nodule $\geq$ 3 mm) is annotated by up to 4 human annotators.
@@
 Eight biomarkers that are informative in clinical practices are also collected in LIDC: subtlety, calcification, sphericity, margin, lobulation, spiculation, texture, and diameter.
-For clinical descriptions of biomarkers, refer to [45].
+For clinical descriptions of biomarkers, refer to (Opulencia et al., 2011).
 After preprocessing, 854 nodule crops of size, corresponding biomarkers, and labels (benign or malignant) were saved and used for downstream procedures.
@@
 4 Results
-For our experiments using the ImageNet [19], we have been using the ViT-S/8 architecture trained with DINO as a backbone [9], which is available as a pre-trained model in their repository [^1] https://github.com/facebookresearch/dino.
-We first discuss the results of applying CDAM qualitatively and then perform a quantitative comparison with several baseline methods.
+For our experiments using the ImageNet (Deng et al., 2009), we use the ViT-S/8 architecture (with a patch size of $8\times 8$) trained with DINO as a backbone (Caron et al., 2021).
+Instead of using a ViT model that is pre-trained in a self-supervised manner, we demonstrate CDAMs directly from an alternative ViT architecture trained with supervised weakly through hashtags (SWAG) (Singh et al., 2022) in Section A.5.
+We further demonstrate applicability of CDAM on computed tomography (CT) scans, using DINO (Caron et al., 2021) in Section 5, as well as using Data-efficient image Transformers (DeIT) (Touvron et al., 2021) and DINOv2 (Oquab et al., 2024) in Section A.5.2.
 4.1 Qualitative Evaluation
 4.1.1 CDAM for a known class
-To demonstrate our method, we have trained a linear classifier on the ImageNet [19] that achieved an accuracy of $77.0\%$.
+Figure 3: Visual comparison of different importance estimators which obtain explanations w.r.t.
+different output classes.
+Similar heat maps regardless of target classes imply that importance estimators are not class discriminative.
+Attention maps (AM) are, by design, not class discriminative.
+Orange and blue colors correspond to positive and negative values, respectively.
+To demonstrate our method, we have trained a linear classifier on the ImageNet (Deng et al., 2009) that achieved an accuracy of $77.0\%$.
 We use random resized cropping and horizontal flipping with PyTorch default arguments as augmentation, Adam optimizer with learning rate $3\times 10^{-4}$, batch size of 128, and train for 10 epochs.
@@
 Note that CDAM does not assign non-zero $S_{i,c}$ to tokens that have zero attention in attention maps, see also the scatter plot between CDAM and attention map Fig.
-11.
-CDAM therefore appears to compute $S_{i,c}$ by multiplying the attention with the relevance of the corresponding token for the targeted class (Section A.1).
-This is a nice feature since CDAM inherits the high-quality object segmentations present in attention maps.
+S1.
+CDAM therefore appears to compute $S_{i,c}$ by multiplying the attention with the relevance of the corresponding token for the targeted class.
+We discuss the proportionality of CDAM to the attention scores in Section A.1.
+It is a nice feature since CDAM inherits the high-quality object segmentations present in attention maps.
 This qualitative evaluation of shrinkage and sparsity of CDAM is reflected in the quantitative measures for compactness (Section 4.2.1).
@@
 In Fig.
-6, targeting the burger or the hotdog does not seem to change importance scores from partial LRP, Input $\times$ Grad, and IntGrad.
+3, targeting the burger or the hotdog does not seem to change importance scores from partial LRP, Input $\times$ Grad, and IntGrad.
 Relevance propagation demonstrates some selective focus on the ROI, but is not as discriminative as CDAM (Fig.
-6).
+3).
 This visual impression of class-sensitivity is supported by the quantitative evaluation in Section 4.2.2.
-Additional examples of CDAM and other interpretability methods are shown in the Appendix A.
-Smooth CDAM appears to reduce the regularization of importance scores, highlighting, for instance, a larger portion of the truck and cars in the background in Fig.
-5.
-This is quite different from how SmoothGrad behaves for CNNs [15], where it usually denoises the importance scores leading to a stronger focus on the ROI.
-On the other hand, Integrated CDAM can’t, for most samples, be visually distinguished from vanilla CADM even though its performance in the benchmarks is quite different (below).
+Additional examples of CDAM and other interpretability methods are shown in Section A.3.
 4.1.2 CDAM for a latent concept
-Concept embeddings $l_{c}$ have been obtained by averaging 30 latent representations of images depicting the concept of interest.
-These 30 images were randomly selected from the corresponding classes in the ImageNet validation set [19] and the CDAM in Fig.
+Figure 4: Class-Discriminative Attention Maps (CDAM) for user-defined concepts (Section 3.2).
+The concept embedding $l_{c}$ has been obtained by averaging the latent representations of 30 images that include the common concept.
+In each instance, from left to right: Sample image, AM, and CDAM.
+Orange and blue correspond to positive and negative values, respectively.
+Concept embeddings $l_{c}$ have been obtained by averaging latent representations of 30 images.
+Those images were randomly selected from a class in the ImageNet validation set (Deng et al., 2009).
+Then, CDAMs in Fig.
 4 were obtained using Eq.
 4.
-We observe that CDAM reveals that the model clearly distinguishes the parts of the images that are relevant for the chosen concept from the rest of the image.
-For example, for the concept hammer (bottom right), CDAM primarily highlights the hammer.
-CDAM also provides positive importance scores for the wooden handles of the screwdrivers and the top of the ratchet, which are indeed relevant for this concept.
+We observe that CDAM reveals that the model clearly distinguishes the parts of the images that are relevant for the latent concepts shared by selected images from the rest of the image.
+For example, when having selected images containing hammer (bottom right), CDAM primarily highlights the hammer.
+While the annotated classes were used for this experiment, the concepts shared by those images are likely more nuanced, specific, and complex than the class itself.
+For example, CDAM also provides positive importance scores for the wooden handles of the screwdrivers and the socket attached the ratchet, which imply relevance for the concepts shared by 30 images selected from the class hammer.
+Semantically, 30 images that are selected from a class hammer have a wooden handle (i.e., concept), which is shared by multiple tools.
+Similarly, the socket at the top of the ratchet shares similarities with typical metal heads at the top of the hammer.
 Unrelated tools or parts appear in blue indicating that the model deems them to be unrelated to the concept hammer.
-Just as CDAM is sensitive to the chosen class, it also produces attention maps that are sensitive to the chosen concept (Fig.
-4, right).
+Figure 5: When targeting different classes (left) or concepts (right), the resulting CDAMs are distinct and align with the corresponding objects.
+Note that green parts of vegetables are shown to activate the zuccini class (left).
+The wrinkles in the elephant’s trunk appear to be related to the zebra concept, semantically mistaken by the model as zebra stripes.
+CDAMs are discriminative with respect to the target concepts shared by selected samples (Fig.
+5, right).
 This makes the proposed method particularly valuable for understanding the representations the ViT has learned, since the explanations for different concepts can give a complementary view.
-For instance, CDAM distinguishes african elephant and zebra.
-However, CDAM also shows that parts of the elephant’s trunk are counter-evidence for the concept african elephant (Fig.
-4, right) which is initially counter-intuitive.
-However, when targeting the concept zebra, the model deems this part of the trunk to be similar to a zebra.
-Speculatively, the trunk may be too similar to the zebra’s muzzle or the trunk’s wrinkles may be mistaken for the zebra’s stripes.
+For example, by selecting 30 images that contain african elephant, CDAM highlights areas that are important for concepts which coincide largely with the body of an elephant.
+It also shows that parts of the elephant’s trunk are counter-evidence for the concepts shared by images with african elephant (Fig.
+5, right).
+For the same sample image, we can also select 30 images with zebra, which suggests that the model deems the part of the elephant’s trunk to be similar to concepts shared by those images.
+Although we selected a group of images based on annotated classes, which are denoted in Fig.
+5 (right), to illustrate concept-based CDAMs, generally the concepts shared by selected images are not identical to the classes themselves.
+One could manually provide a set of images with a shared concept that is not originally annotated in the dataset.
+For example, 10 images that contain zebras may share semantic concepts (e.g., “stripes” and “tail”), that make up the class zebra.
+We demonstrate this by selecting a set of 10 images, outside of the ImageNet, that exclusively show stripes and another set of 10 images that contain diverse animal tails.
+Shown in Fig.
+6, CDAM for stripes highlights the body of the zebra, but not the body of the cheetah.
+On the other hand, CDAM for tails tends to focus on the tail of the cheetah.
+Note that 10 images of diverse animal tails include body parts of animals, often rear ends and legs.
+Figure 6: Concept-based CDAM based on 10 images that show stripes or contain animal tails.
 4.2 Quantitative evaluation
@@
 4.2.1 Evaluation of correctness
+Fidelity Box Sensitivity
+[…]
+CDAM 404 1038 632 12.2
+Integrated CDAM 375 1045 669 13.0
+Smooth CDAM 282 1110 827 12.4
+Relevance Propagation 315 957 643 10.0
+Partial LRP 489 837 348 4.5
+Attention map 452 846 393 4.8
+SmoothGrad 534 529 -4 0.4
+IntGrad 226 802 574 7.6
+GradCAM 300 944 643 8.6
+Input $\times$ Grad 273 713 439 3.5
+Table 1: Evaluation of correctness by fidelity and box sensitivity.
+In $X_{\text{multi}}$, importance scores for the target class were used to perturb pixels.
+The changes in the model output w.r.t.
+the perturbation percentage ($f_{\text{LIF-MIF}}$; Fig.
+S2) or the box size ($f_{\text{box}}$; Fig.
+S3) are quantified by the area under the curves.
 […]
@@
 4.2.2 Evaluation of class discrimination
+Fidelity Box Sensitivity
+[…]
+CDAM 739 14.6
+Integrated CDAM 708 14.8
+Smooth CDAM 823 13.4
+Relevance Propagation 293 6.6
+Partial LRP 3 0.0
+Attention map 0 -0.2
+SmoothGrad -11 -0.1
+IntGrad 540 7.6
+GradCAM 681 9.4
+Input $\times$ Grad 433 3.4
+Table 2: Evaluation of class discrimination by differences in fidelity and box sensitivity metrics.
+By using the $X_{\text{multi}}$ dataset, we measure how sensitive importance scores are to changing their target class (Details in Section 3.5).
+Area under the $\Delta f_{\text{LIF-MIF}}$ (Fig.
+S4) and $\Delta f_{\text{box}}$ curves (Fig.
+S5).
 […]
@@
 The importance scores for CDAM become anti-correlated with the model output when the wrong importance scores are used for the perturbation (Fig.
-10).
+S5).
 This anti-correlation probably results from the fact that pixels that are evidence for the correct target class are counter-evidence for the wrong class.
 This result corroborates the visual impression that CDAM correctly assigns importance scores with opposite signs to the objects corresponding to the targeted class and other objects that are present (Fig.
-6).
+3).
 This is also reflected in the negative values of $f_{\text{LIF}^{\prime}}-f_{\text{MIF}^{\prime}}$ for CDAM (Fig.
-9), indicating that the ranking from least to most important has been, at least partially, reversed.
+S4), indicating that the ranking from least to most important has been, at least partially, reversed.
 4.2.3 Evaluation of compactness
-Sparsity and shrinkage are evaluated by our compactness evaluation, which is one of the major desired properties in explainability of deep learning [44].
-While sparsity may not always imply the most accurate estimation at a single data point, a bias-variance trade-off is well-known in machine learning [47].
+[…]
+CDAM 0.88
+Integrated CDAM 0.88
+Smooth CDAM 0.75
+Relevance Propagation 0.48
+Partial LRP 0.74
+Attention map 0.56
+SmoothGrad 0.8
+IntGrad 0.84
+GradCAM 0.54
+Input $\times$ Grad 0.88
+Table 3: Compactness, measuring sparsity and shrinkage, is defined by the fraction of pixels with an importance score lower or equal to 5% of the maximum score, evaluated on the $X_{\text{multi}}$ dataset.
+Sparsity and shrinkage are evaluated by our compactness evaluation, which is one of the major desired properties in the explainability of deep learning (Nauta et al., 2023a).
+While sparsity may not always imply the most accurate estimation at a single data point, a bias-variance trade-off is well-known in machine learning (Lahlou Kitane, 2022).
 Ideally, compact explanations (e.g., sparse and shrunken importance estimators) are preferred, when other properties such as correctness and class sensitivity are constant.
@@
 5 Application to Medical Images
-To demonstrate CDAM for another use case we apply it to lung nodule malignancy and biomarker prediction using the LIDC dataset [23].
+To demonstrate CDAM for another use case we apply it to nodule malignancy and biomarker prediction using the Lung CT scans in the LIDC dataset (Armato III et al., 2011).
+Here, we show CDAMs from DINO-based models, we also showcase use of DeIT (Dosovitskiy et al., 2020; Touvron et al., 2021) and DINOv2 (Oquab et al., 2024; Darcet et al., 2024) backbones that are fine-tuned on the LIDC dataset in Section A.5.2.
 5.1 Malignancy Prediction
@@
 Training, validation, and test sets (in the ratios of 0.7225, 0.1275, 0.15) were stratified by and balanced according to these labels, e.g., benign and malignant.
-We fine-tuned a ViT model, with a DINO backbone pretrained on the ImageNet, for 50 epochs.
+We fine-tuned a ViT model, with a DINO backbone pre-trained on the ImageNet, for 50 epochs.
 In a parameter sweep, we varied the number of trainable layers ($10-50$) and dropout rates ($0.0-0.09$), where the learning rate was exponentially decaying ($\alpha=0.0003$ and $\beta=0.95$).
-The best accuracy on the test set of 0.85 was obtained with 50 trainable layers and dropout rate of 0.003222It might be possible to increase the model performance via further investigation.
+The best accuracy on the test set of 0.85 was obtained with 50 trainable layers and dropout rate of 0.003111It might be possible to increase the model performance via further investigation.
 Here, we use a straightforward model and training scheme to demonstrate the application of CDAM on medical images..
@@
 The obtained attention maps suggest that the model focuses on patches with nodule fragments (Fig.
-12(b)).
+S11(b)).
 However, we note that attention maps almost always focus on pulmonary nodules without taking into account the downstream classification into benign and malignant samples.
@@
 We turn our attention to clinical biomarkers in LIDC that are routinely used by medical practitioners.
-Inspired by the concept bottleneck model (CBM) [48], we fine-tuned a ViT model pre-trained with DINO on 8 biomarkers (subtlety, calcification, sphericity, margin, lobulation, spiculation, texture, and diameter).
+Inspired by the concept bottleneck model (CBM) (Koh et al., 2020), we fine-tuned a ViT model pre-trained with DINO on 8 biomarkers (subtlety, calcification, sphericity, margin, lobulation, spiculation, texture, and diameter).
 Essentially, a regression model was built on a pre-trained ViT model with the mean squared error (MSE) as loss function.
-For investigation of LIDC and incorporation of CBM, see [49].
+For investigation of LIDC and incorporation of CBM, see (Brocki & Chung, 2023a).
 CDAMs were obtained for 8 biomarkers (Fig.
-13).
+7).
 In the context of the margin, patches including the edge of a nodule get positive CDAM scores, whereas patches corresponding to the nodule body get negative CDAM scores.
@@
 In a majority of cases, CDAM highlights spicules with positive values, whereas the inner bodies of nodules (i.e., not spicules) tend to get low importance scores.
+Figure 7: Explanations for the biomarker prediction model.
+Eight biomarkers available in the LIDC dataset are used to fine-tune a ViT regression model with a DINO backbone.
+CDAMs provide distinct and relevant explanations for different biomarkers.
+For more examples, see Section A.4.
 6 Discussion
 Class-Discriminative Attention Map (CDAM) is a gradient-based extension of attention maps, that provides strong discrimination with respect to a chosen class or concept, while exhibiting excellent fidelity and compactness.
-We introduce additional variants of CDAM, inspired by SmoothGrad ([15]) and IntGrad [16]), which improve the performance on the quantitative benchmarks.
-Our proposed methods retain appealing characteristics of attention maps, namely high-quality semantic segmentation in self-supervised ViTs.
+Our proposed methods retain appealing characteristics of attention maps, namely their high-quality semantic segmentation.
 The proposed importance estimation for CDAM scales the attention scores by how relevant they are for a target class or concept (Section A.1).
 Therefore, zero attention scores remain zero in CDAM and it produces compact explanations with high sparsity and shrinkage.
+Besides explaining predictions of the classifier on the top of ViT, CDAM can also provide importance scores specific to the user-defined concept.
+In the context of self-supervised models, where class labels are absent, this makes it a valuable technique to investigate the latent representations learned by the model.
 Post-hoc explanation methods for DNNs aim to make the full decision-making process of the model more transparent by providing an approximation with certain desirable properties, such as correctness, sensitivity, and compactness.
-Particularly, many explanation methods that estimate feature importance are not very useful because they give almost identical results when targeting different classes in the model output [27, 38].
-Even randomization in the model weights is often shown to have minimal impact on explanations (i.e., saliency maps [38]).
+Particularly, many explanation methods that estimate feature importance are not very useful because they give almost identical results when targeting different classes in the model output (Rudin, 2019; Adebayo et al., 2018).
+Even randomization in the model weights is often shown to have minimal impact on explanations (i.e., saliency maps (Adebayo et al., 2018)).
 We find, both qualitatively and quantitatively, that CDAM is highly sensitive to the targeted class, assigning positive importance scores to objects corresponding to the target class and negative ones to other (semantically distinct) objects in the image.
-Besides explaining predictions of the classifier on the top of ViT, we introduce several variants.
-Firstly, by probing the latent space of the ViT (Fig.
-4), self-supervised CDAM provides importance scores specific to the user-defined concept.
-In the context of SSL, where class labels are absent, this makes it a valuable technique to investigate the representations learned by the model, as shared by user-selected samples.
-To some degree, this can alternatively be done by looking at the attention maps of the different heads, which tend to specialize in different categories of objects or parts thereof [9].
-This approach is, however, limited to the categories that the attention heads happened to have learned.
-Secondly, we introduce Smooth and Integrated CDAM, which essentially average a series of (vanilla) CDAMs.
-Notably, instead of adding noise to or modifying the input images as in SmoothGrad [15]) or IntGrad [16], respectively, our methods act on tokens in the final transformer block.
-Our application on medical images examplifies the need for fine-grained class-sensitive explanations offered by CDAM.
-Other explainability methods including attention maps are often highlighting whole nodules and tumors.
-But they are uninformative about the innerworkings of the ViT model predicting malignancy or biomarkers.
-While medical implication of these explanations are beyond the scope of this paper, the proof of concept application demonstrates usability of CDAM in cancer images.
-As the methods of explainable AI (XAI) becomes incorporated into practices, we foresee a greater level of human-computer interaction where class sensitivity will play a critical role for utility of explanations.
-Understanding the decision-making process of deep neural networks is intrinsically challenging due to the vast parameter space and non-linearity, which lead to complex interactions among the patch tokens in ViT.
-Post-hoc explanation methods for DNNs approximate the full operating characteristics of the model with certain desirable properties.
-Our results suggest that CDAM is capable of achieving a high degree of correctness, contrastivity, and compactness.
-Note that CDAM is not explicitly designed to satisfy these criteria.
-For example, we could impose further sparsity constraints that would increase compactness.
-However, over-optimizing an explanation method for certain properties could have undesirable effects [^3] Goodhart’s Law states that “when a measure becomes a target, it ceases to be a good measure” [50]..
-While this paper focuses on a ViT model, CDAM is applicable more broadly to all transformer models with self-attention, including large language models (LLMs) [51].
-Acknowledgments
-This work was funded by the ERA-Net CHIST-ERA grant [CHIST-ERA-19-XAI-007] long-term challenges in ICT project INFORM (ID: 93603), by the National Science Centre (NCN) of Poland [2020/02/Y/ST6/00071].
-This research was carried out with the support of the Interdisciplinary Centre for Mathematical and Computational Modelling University of Warsaw (ICM UW) under computational allocation no GDM-3540; the NVIDIA Corporation’s GPU grant; and the Google Cloud Research Innovators program.
+We introduce Smooth and Integrated CDAM, which essentially average a series of (vanilla) CDAMs.
+Notably, instead of adding noise to or modifying the input images as in SmoothGrad (Smilkov et al., 2017) or IntGrad (Sundararajan et al., 2017), respectively, our methods act on tokens in the final transformer block.
+While we conducted a number of quantitative and qualitative evaluations, there is no single XAI method that triumph in all scenarios.
+However, considering that Smooth and Integrated CDAM require computing $n$ vanilla CDAMs (e.g., $n=50$), we recommend first trying vanilla CDAMs.
+If high fidelity is singularly desirable, we suggest utilizing Smooth CDAM.
+Our application on medical images exemplifies the need for fine-grained class-sensitive explanations offered by CDAM.
+Other explainability methods including attention maps often highlight whole nodules and tumors.
+However, they are uninformative about the inner workings of the ViT model predicting malignancy or biomarkers.
+As the methods of explainable AI (XAI) are considered for practical implementations, the investigation of human interactions with different explainability methods represents a promising direction for future studies.
+Broader Impact Statement
+The introduction of our importance estimators would enhance the transparency, trustworthiness and accountability of vision transformer models.
+By providing clearer insights into model decision-making processes, accurate explainability methods help foster broader adoption of AI technologies.
+However, the explanations could be manipulated by adversarial input perturbation or data poisoning.
+In turn, enhanced explanations may inadvertently expose vulnerabilities, making models more susceptible to adversarial attacks or malicious exploitation.
+Therefore, it is imperative to continuously monitor and ensure the robustness of explanation methods while implementing safeguards to mitigate the risks associated with adversarial attacks.
 Appendix A Appendix
@@
 Empirically, in all of our evaluation tasks and applications (e.g., Fig.
-11), we observe that features with attention scores $A_{i}=\text{softmax}\left(\frac{Q_{\texttt{[CLS]}}K^{\top}_{i}}{\sqrt{d_{k}}}\right)$ of 0 have class-discriminative attention scores of 0:
+S1), we observe that features with attention scores $A_{i}=\text{softmax}\left(\frac{Q_{\texttt{[CLS]}}K^{\top}_{i}}{\sqrt{d_{k}}}\right)$ of 0 have class-discriminative attention scores of 0:
 $A_{i}=0\Rightarrow S_{i,c}=0.$ (9)
@@
 To better understand this observation, we first consider a concept-based CDAM with a single attention head. Eq. 4 is expanded:
-$S_{i,c}$ $=\sum_{j}T_{ij}\frac{\partial g(l,l_{c})}{\partial T_{ij}}$
-$=\sum_{j}T_{ij}\frac{\partial}{\partial T_{ij}}\sum_{k}l_{k}l_{c,k}$
-$=\sum_{j}T_{ij}\frac{\partial}{\partial T_{ij}}\sum_{k}h_{k}\left(\sum_{m}A_{m}V_{m}+\texttt{[CLS]}\right)l_{c,k}$
-$=\sum_{j}T_{ij}\sum_{k,m,n}\nabla_{n}h_{k}\times\left(\frac{\partial A_{m}}{\partial T_{ij}}V_{mn}+A_{m}\frac{\partial V_{mn}}{\partial T_{ij}}\right)l_{c,k}$
-$=\sum_{j,k,m,n}T_{ij}\nabla_{n}h_{k}\frac{\partial A_{m}}{\partial T_{ij}}V_{mn}l_{c,k}+\sum_{k,n}A_{i}V_{in}\nabla_{n}h_{k}l_{c,k},$ (10)
+$S_{i,c}$ $=\sum_{j}T_{ij}\frac{\partial g(l,l_{c})}{\partial T_{ij}}$ (10)
+$=\sum_{j}T_{ij}\frac{\partial}{\partial T_{ij}}\sum_{k}l_{k}l_{c,k}$ (11)
+$=\sum_{j}T_{ij}\frac{\partial}{\partial T_{ij}}\sum_{k}h_{k}\left(\sum_{m}A_{m}V_{m}+\texttt{[CLS]}\right)l_{c,k}$ (12)
+$=\sum_{j}T_{ij}\sum_{k,m,n}\nabla^{n}h_{k}\times\left(\frac{\partial A_{m}}{\partial T_{ij}}V_{mn}+A_{m}\frac{\partial V_{mn}}{\partial T_{ij}}\right)l_{c,k}$ (13)
+$=\sum_{j,k,m,n}T_{ij}\nabla^{n}h_{k}\frac{\partial A_{m}}{\partial T_{ij}}V_{mn}l_{c,k}+\sum_{k,n}A_{i}V_{in}\nabla^{n}h_{k}l_{c,k},$ (14)
 […]
-Eq. 9 implies that in Section A.1, all terms that are not proportional to $A_{i}$ are zero or cancelled out and we are left with
+Eq. 9 implies that in Eq. 10, all terms that are not proportional to $A_{i}$ are zero or cancelled out and we are left with
 $S_{i,c}=A_{i}V_{i}\cdot\nabla(h\cdot l_{c}).$ (15)
@@
 In the special case that $h$ is the identity function, we obtain an even simpler interpretation of CDAM due to $\nabla_{n}h_{k}(f)=\partial_{f_{n}}f_{k}=\delta_{nk}$.
-Assuming $\frac{\partial A_{m}}{\partial T_{ij}}=0$, Section A.1 simplifies to $S_{i,c}=A_{i}V_{i}\cdot l_{c}$, which means that $A_{i}$ is scaled by the similarity of $V_{i}$ and $l_{c}$.
+Assuming $\frac{\partial A_{m}}{\partial T_{ij}}=0$, Eq.
+10 simplifies to $S_{i,c}=A_{i}V_{i}\cdot l_{c}$, which means that $A_{i}$ is scaled by the similarity of $V_{i}$ and $l_{c}$.
 We note that introducing $h$ attention heads does not change the above derivations qualitatively.
-Instead of the sum $Z=\sum_{m}A_{m}V_{m}$, there will be a term $\text{Concat}(Z^{1},Z^{2},..,Z^{h})W^{O},W^{O}\in\mathbb{R}^{hd_{v}\times d_{\text{model}}}$ so that in the case of multi-head attention the equivalent of Section A.1 reads
+Instead of the sum $Z=\sum_{m}A_{m}V_{m}$, there will be a term $\text{Concat}(Z^{1},Z^{2},..,Z^{h})W^{O},W^{O}\in\mathbb{R}^{hd_{v}\times d_{\text{model}}}$ so that in the case of multi-head attention the equivalent of Eq.
+10 reads
 $S_{i,c}=\text{Concat}\left(Y^{1},Y^{2},..,Y^{h}\right)W^{O}\nabla(h\cdot l_{c}),\;Y^{n}=T_{i}\sum_{m}\frac{\partial A^{n}_{m}}{\partial T_{i}}V^{n}_{m}+A^{n}_{i}V^{n}_{i}$ (16)
@@
 9 to hold all terms that are not proportional to $A_{i}$ either vanish or cancel out.
-For multi-head attention we thus arrive at
+For multi-head attention, we thus arrive at
 $S_{i,c}=\text{Concat}\left(Y^{\prime 1},Y^{\prime 2},..,Y^{\prime h}\right)W^{O}\nabla(h\cdot l_{c}),\;Y^{\prime n}=A^{n}_{i}V^{n}_{i},$ (17)
@@
 Each of the summands is proportional to the corresponding $A^{n}$, and the interpretation of CDAM from the single-head case therefore extends to the multi-head one.
-Although our observation $A_{i}=0\Rightarrow S_{i,c}=0$ implies that all terms that are not proportional to $A_{i}$ in Section A.1 should vanish, we don’t have a theoretical argument why that would be the case.
+Although our observation $A_{i}=0\Rightarrow S_{i,c}=0$ implies that all terms that are not proportional to $A_{i}$ in Eq.
+10 should vanish, we don’t have a theoretical argument why that would be the case.
 It seems unlikely that the various terms in the first sum cancel out, which leaves $\frac{\partial A_{m}}{\partial T_{ij}}=0$ as an explanation.
@@
 Analogously to the above single-head discussion we have
-$S_{i,c}$ $=\sum_{j}T_{ij}\frac{\partial f_{c}}{\partial T_{ij}}$
-$=\sum_{j}T_{ij}\frac{\partial}{\partial T_{ij}}\sum_{k}l_{k}W^{C}_{kc}$
-$=\sum_{j,k,m,n}T_{ij}\nabla_{n}h_{k}\frac{\partial A_{m}}{\partial T_{ij}}V_{mn}W^{C}_{kc}+\sum_{k,n}A_{i}V_{in}\nabla_{n}h_{k}W^{C}_{kc},$ (14)
+$S_{i,c}$ $=\sum_{j}T_{ij}\frac{\partial f_{c}}{\partial T_{ij}}$ (18)
+$=\sum_{j}T_{ij}\frac{\partial}{\partial T_{ij}}\sum_{k}l_{k}W^{C}_{kc}$ (19)
+$=\sum_{j,k,m,n}T_{ij}\nabla^{n}h_{k}\frac{\partial A_{m}}{\partial T_{ij}}V_{mn}W^{C}_{kc}+\sum_{k,n}A_{i}V_{in}\nabla^{n}h_{k}W^{C}_{kc},$ (20)
 and the only difference is that the directional derivative now acts on $\sum_{k}h_{k}W_{kc}$, i.e.
@@
 All the intuitions we have gained from the concept-based CDAM therefore extend to the class-based one.
-A.2 Results for a random subset of ImageNet
-Figure S1: Box sensitivity curves obtained from the random ImageNet subset
-Figure S2: MIF/LIF perturbation curves obtained from the random ImageNet subset.
+A.2 Quantitative Evaluations
+Figure S1: Scatter plot of raw attention scores and token importance scores obtained from 200 randomly selected samples of the ImageNet validation set $X_{\text{val}}$.
+Figure S2: Fidelity of importance estimators based on MIF/LIF perturbation curves obtained from the $X_{\text{multi}}$ dataset.
+The change in the model output is measured as a proportion of input features are perturbed according to the Least Important First (LIF) or the Most Important First (MIF) orders.
+Importance estimators are evaluated by the area between LIF and MIF perturbation curves, which is equivalent to the area underneath $f_{\text{LIF}}$-$f_{\text{MIF}}$ (right).
+See Table 3.
+Figure S3: Correctness of importance estimators as measured by box sensitivity evaluation.
+The change in the model output (logit) is measured as pixels within a $s\times s$ box are perturbed.
+The sum of their importance scores would correlate with the change in the model output, if importance estimators are accurate.
+This is repeated 100 times per sample and the Pearson correlation between the sum of importance scores and the change in the model output is calculated, based on the $X_{\text{multi}}$ dataset.
+Figure S4: MIF/LIF perturbation curves for the evaluation of class-discrimination.
+Different classes are targeted for obtaining importance scores and measuring model output (Section 3.5), a large $\Delta A_{\text{box}}$, therefore, indicates importance scores that are class-discriminative.
+See Table 2
+Figure S5: Box sensitivity curves for the evaluation of class-sensitivity
+Figure S6: Box sensitivity curves obtained from the random ImageNet subset $X_{\text{val}}$. Refer to Section 3.5 for data description.
+Figure S7: MIF/LIF perturbation curves obtained from the random ImageNet subset $X_{\text{val}}$. Refer to Section 3.5 for data description.
 […]
@@
 IntGrad 411 1055 644 8.3
+GradCAM 515 1249 733 9.0
 Input x Grad 470 935 465 3.1
 Table S1: Area under the curve for MIF, LIF and LIF-MIF curves (Fig. S7) and box sensitivity curves (Fig. S6) obtained from the random ImageNet subset.
-A.3 Examples of CDAM for a known class on the ImageNet
-Figure S3: Examples for CDAM and other explanation methods on the ImageNet. The indicated class is the one predicted by the model and is the target for the explanation methods.
-Figure S4: Examples for CDAM and other explanation methods on the ImageNet. The indicated class is the one predicted by the model and is the target for the explanation methods.
-Figure S5: Examples for CDAM and other explanation methods on the ImageNet. The indicated class is the one predicted by the model and is the target for the explanation methods.
-A.4 Examples for CDAM on the Lung Image Database Consortium (LIDC)
-Figure S6: Examples for the Attention Map and CDAM based on a ViT malignancy model trained on the lung nodule data from LIDC.
+A.3 Examples of CDAMs from the ImageNet
+Figure S8: Additional examples for CDAM and other explanation methods on the ImageNet. The indicated class is the one predicted by the model and is the target for the explanation methods.
+Figure S9: Additional examples for CDAM and other explanation methods on the ImageNet. The indicated class is the one predicted by the model and is the target for the explanation methods.
+Figure S10: Additional examples for CDAM and other explanation methods on the ImageNet. The indicated class is the one predicted by the model and is the target for the explanation methods.
+A.4 Examples for CDAMs from lung CT scans
+Figure S11: Explanations for the malignancy prediction model.
+A ViT-based classifier fine-tuned on the lung CT scans (LIDC) was used to predict benign vs.
+malignant lung nodules, followed by obtaining their attention maps and CDAMs.
+Orange and blue correspond to positive and negative values, respectively.
 Figure S12: Examples for the Attention Map and CDAM based on a ViT biomarker model trained on the lung nodule data from LIDC.
 Figure S13: Examples for the Attention Map and CDAM based on a ViT biomarker model trained on the lung nodule data from LIDC.
+A.5 Additional ViT architectures and training strategies
+While the main manuscript has used ViT models pre-trained using DINO (Caron et al., 2021), other architectures and pre-trained weights can be used for CDAM.
+In this section, we show examples of alternative ViT models that are trained in alternative manners.
+As expected, pre-trained ViT models that produce high-quality attention maps result in high-quality CDAMs.
+In particular, note that the patch size of DINO (Caron et al., 2021) is $8\times 8$ compared to much larger patch sizes used in other backbones.
+Table S2: Pre-trained ViT models and training strategies.
+Name Model sizes Patch sizes References
+DeIT (original ViT) 86M 16 Dosovitskiy et al. (2020); Touvron et al. (2021)
+SWAG 86M 16 Singh et al. (2022)
+DINOv2 (with registers) 21M, 86M 14 Oquab et al. (2024); Darcet et al. (2024)
+DINO (used in the main text) 21M, 85M 8, 16 Caron et al. (2021)
+A.5.1 Alternative ViT models for the ImageNet
+We use a ViT model pre-trained on the ImageNet-1K dataset using supervised weakly through hashtags (SWAG) (Singh et al., 2022).
+Unlike DINO and other self-supervised ViT models, the SWAG model has been trained end-to-end.
+Therefore, without pre-training, we can obtain attention maps and CDAMs directly.
+Three examples are shown in Fig.
+S14.
+Figure S14: Attention maps and CDAMs from the ImageNet-1K dataset, based on the pre-trained ViT model backbone using SWAG Singh et al. (2022).
+In general, ViT models require a massive amount of training data and therefore often trained in self-supervised manner.
+As shown in DINO (Caron et al., 2021), self-supervised learning can not only improve performance of fine-tuned models, but also improve attention maps.
+CDAM relies on attention maps, therefore the quality of attention maps strongly influence that of CDAM.
+Note that the SWAG model has a patch size of $16\times 16$, which results in lower-resolution attention maps and CDAMs.
+A.5.2 Alternative ViT models for the LIDC
+We demonstrate the use of CDAM with ViT models trained using Data-efficient image Transformers (DeIT) (Touvron et al., 2021) and DINOv2 (Oquab et al., 2024).
+Note that despite the use of the same acronym, DINOv2 is substantially different from DINO (Caron et al., 2021) due to using a different training strategy introduced in iBOT (Zhou et al., 2022) on a much larger scale (Oquab et al., 2024).
+The original DINO (Caron et al., 2021) was pre-trained on the ImageNet, whereas DINOv2 (Oquab et al., 2024) used a newly curated dataset LVD-142M.
+Additional tokens called “registers” were introduced to improve attention maps of DINOv2 (Darcet et al., 2024).
+The LIDC dataset was split into 5 folds and stratified according to malignancy status.
+The best-performing models were chosen based on means of Accuracy (ACC) and Mean Squared Error (MSE) over 5 folds, in the validation set.
+In a parameter sweep, we varied the number of trainable layers (10 - all), dropout rates in the backbone (0.0 - 0.12), batch size (8 - 32) and learning rate scheduler parameters.
+The learning rate was scheduled with CyclicLR scheduler.
+Binary Cross Entropy Loss and Huber Loss were used to train classification and regression models, respectively.
+Weights optimization was performed with Adam optimizer and random rotation was applied as data augmentation.
+For malignancy classification, the best-performing models based on DeIT and DINOv2 achieved a mean ACC of 0.896 and 0.904, respectively.
+For biomarker regression, the best-performing models based on DeIT and DINOv2 archived the mean MSE of 0.409 and 0.35, respectively.
+Figure S15: Attention Maps and CDAMs for malignancy models using ViT backbones pre-trained with DeIT and DINOv2.
+Figure S16: Attention Maps and CDAMs for biomarkers models using ViT backbones pre-trained with DeIT
+Figure S17: Attention Maps and CDAMs for biomarkers models using ViT backbones pre-trained with DINOv2